### v1

# Slant sums of quiver gauge theories

Abstract
We define a notion of slant sum of quiver gauge theories, a type of surgery on the underlying quiver. Under some mild assumptions, we relate torus fixed points on the corresponding Higgs branches, which are Nakajima quiver varieties. Then we prove a formula relating the quasimap vertex functions before and after a slant sum, which is a type of “branching rule” for vertex functions.
Our construction is motivated by a conjecture, which we make here, for the factorization of the vertex functions of zero-dimensional quiver varieties. The branching rule allows this conjecture to be approached inductively. In special cases, it also provides a formula for the $\hbar=q$ specialization of vertex functions for quiver varieties not necessarily of Dynkin type as a sum over reverse plane partitions. When passed through the quantum Hikita conjecture, such expressions provide conjectural formulas for graded traces of Verma modules on the 3d mirror dual side.
We also consider the Coulomb side. We make some conjectures reflecting what can be seen on the Higgs side and prove them in ADE type. We study slant sums of Coulomb branches and their quantizations. We prove that for one-dimensional framing, slant sum operation on the Coulomb branch side corresponds to taking products.
1 Introduction
1.1 Slant sums of heaps
The slant sum of two posets was defined by Proctor in [38] to break certain posets into simpler, slant irreducible pieces. Some particular posets to which this applies are those that appear as heaps $H(w)$ of fully commutative elements $w$ of a Weyl group $W$, see [39]. Such posets are defined in a straightforward combinatorial way from a choice of reduced word for $w$.
It was noticed in [18] that the heap of a minuscule element gives rise to a module over the preprojective algebra of the corresponding Dynkin quiver. Equivalently, it gives rise to a point in a Nakajima quiver variety. When $w$ is dominant minuscule, the corresponding quiver variety is a single point, and the preprojective algebra module studied in [18] is a choice of representative of this point.
Suppose $w$, $w^{(1)}$, and $w^{(2)}$ are dominant minuscule elements of some Weyl groups and let $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ be the respective zero-dimensional Nakajima quiver varieties. The claim that the heap of $w$ decomposes as a slant sum of the other two heaps, written $H(w)=H(w^{(1)})\#H(w^{(2)})$, is equivalent to saying that a representative of the unique point in $\mathcal{M}$ can be built from representatives of the unique points in $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$. It is thus reasonable to call $\mathcal{M}$ the slant sum of the two other quiver varieties.
The aim of this paper is to revisit slant sums from the perspective of quiver varieties, or more precisely, quiver gauge theories. If the numerical data used to construct $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$, which now need not be zero-dimensional, is compatible, we define a new quiver variety $\mathcal{M}:=\mathcal{M}^{(1)}\#\mathcal{M}^{(2)}$ which we call the slant sum of the two constituents. Unlike in [18], our procedure quickly leaves the world of Dynkin quivers, which we view as a feature.
1.2 Slant sums of quiver gauge theories
Recall that a quiver gauge theory is specified by a choice of quiver $Q=(Q_{0},Q_{1})$ and $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$. It is convenient to think of $\mathsf{v}_{i}$ as associated to a gauge vertex $i\in Q_{0}$ and $\mathsf{w}_{i}$ as associated to a framing vertex, as in the following picture:
$\mathsf{v}_{1}$$\mathsf{v}_{2}$$\mathsf{v}_{3}$$\mathsf{w}_{1}$$\mathsf{w}_{2}$$\mathsf{w}_{3}$
Let $Q^{(r)}$, $\mathsf{v}^{(r)}$, and $\mathsf{w}^{(r)}$ for $r\in\{1,2\}$ be two such collections. Suppose there exist vertices $\star_{1}\in Q^{(1)}_{0}$ and $\star_{2}\in Q^{(2)}_{0}$ such that $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}$. We say that the two quiver gauge theories are slant summable, and we define their slant sum to be the quiver gauge theory defined by identifying the gauge vertex at $\star_{1}$ with the framing vertex at $\star_{2}$. We will use the symbol $_{\star_{1}}\#_{\star_{2}}$, or just #\#, throughout to denote the effect of this operation on various objects. For example, the quiver for the slant sum of the two quiver gauge theories is $Q=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$. It is obtained by adding a single arrow to the disjoint union of the two constituent quivers.
The local picture is
$\mathsf{v}_{\star_{1}}$$_{\star_{1}}\#_{\star_{2}}$$\mathsf{w}_{\star_{2}}$$\mathsf{w}_{\star_{1}}$$\mathsf{v}_{\star_{2}}$$\mathsf{v}_{\star_{1}}$$\mathsf{v}_{\star_{2}}$$\mathsf{w}_{\star_{1}}$
Let $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ be the associated resolved Higgs branches (equivalently, the Nakajima quiver varieties). From the definitions,
$\dim\mathcal{M}=\dim\mathcal{M}^{(1)}+\dim\mathcal{M}^{(2)}$
and it is tempting to think of #\# as a sort of product which preserves connectedness of the quivers. Actually, one interpretation of Corollary 3.7 is that #\# is closer to being 3d mirror dual to the product, see also Proposition 7.2.
In general, $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ are related by a certain diagram, see (7). But under some additional assumptions, we are able to define a map from certain torus fixed points on $\mathcal{M}^{(1)}\times\mathcal{M}^{(2)}$ to those on $\mathcal{M}$. Let $\mathsf{T}$, $\mathsf{T}^{(1)}$, and $\mathsf{T}^{(2)}$ be the tori acting on the quiver varieties. Let $\mathcal{V}_{i}$ be the tautological bundle on $\mathcal{M}$ for the vertex $i$. Similarly, let $\mathcal{V}^{(1)}_{j}$ and $\mathcal{V}^{(2)}_{k}$ be the tautological bundles on $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$.
Theorem 1.1 (Theorem 2.11).
Suppose that $p^{(1)}\in\left(\mathcal{M}^{(1)}\right)^{\mathsf{T}^{(1)}}$ is split over $\star_{1}$ in the sense of Definition 2.9. There is an embedding [^1] Strictly speaking, this embedding depends mildly on a choice of ordering of some equivariant parameters, see Proposition 2.12.
$\Psi_{p^{(1)}}:\left(\mathcal{M}^{(2)}\right)^{\mathsf{T}^{(2)}}\to\mathcal{M}^{\mathsf{T}}.$
Furthermore, there is an inclusion $\iota:\mathsf{T}\hookrightarrow\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}$ such that $\mathcal{V}_{i}|_{p}=\iota^{*}\mathcal{V}^{(r)}_{i}|_{p^{(r)}}$ for $i\in Q^{(r)}$, $r\in\{1,2\}$.
In the setting of the previous proposition, we will also denote $p^{(1)}\#p^{(2)}:=\Psi_{p^{(1)}}(p^{(2)})$. When $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ are 0-dimensional quiver varieties corresponding to dominant minuscule Weyl group elements and $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}=1$, Theorem 1.1 recovers the classical notion of slant sums of heaps studied in [38], see also [18].
1.3 Slant sums and vertex functions
Our constructions also relate to certain quasimap counts. In the $K$-theoretic enumerative geometry of quiver varieties, the descendant vertex function is a key object [37]. It appears in Okounkov’s enumerative 3d mirror symmetry [1, 36], see also [5, 12, 13]. It is defined as the generating function of equivariant counts of quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$ with a nonsingular condition at $\infty\in\mathbb{P}^{1}$. We denote it by
$V^{(\tau)}_{\mathcal{M}}(z)\in K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathcal{M})[[z]]$
Here, $[[z]]$ stands for power series in a certain cone in variables $z_{i}$ for $i\in Q_{0}$. The descendant $\tau$ is an element of $K_{G_{\mathsf{v}}}(\mathrm{pt})$ where $G_{\mathsf{v}}$ is the gauge group. For a fixed point $p\in\mathcal{M}^{\mathsf{T}}$, we will denote $V^{(\tau)}_{p}:=V^{(\tau)}_{p}(z):=V^{(\tau)}_{\mathcal{M}}(z)|_{p}$.
In the setting of Theorem 1.1, there is a map between the spaces of descendants:
$\iota^{*}:K_{\mathsf{T}^{(1)}\times G_{\mathsf{v}^{(1)}}}(\mathrm{pt})\otimes K_{\mathsf{T}^{(2)}\times G_{\mathsf{v}^{(2)}}}(\mathrm{pt})\to K_{\mathsf{T}\times G_{\mathsf{v}}}(\mathrm{pt}),$ (1)
and it is natural to wonder if there is any relationship between $V^{(\tau)}_{p}$, $V^{(\tau_{1})}_{p^{(1)}}$, and $V^{(\tau_{2})}_{p^{(2)}}$ when $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. This is the content of our main theorem.
To state it, let $\mathsf{QM}_{p^{(1)}}$ denote the moduli space of stable quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}^{(1)}$ which send $\infty$ to $p^{(1)}$. There is an evaluation morphism $\operatorname{ev}_{0}:\mathsf{QM}_{p^{(1)}}\to K_{G_{\mathsf{v}^{(1)}}}(\mathrm{pt})$. For a cocharacter $\sigma:\mathbb{C}^{\times}\to\mathsf{A}^{(2)}\subset\mathsf{T}^{(2)}$ of the framing torus, let $V^{(\tau_{2}),\sigma}_{p^{(2)}}$ be the $\sigma$-twisted vertex function.
Theorem 1.2 (Theorem 3.2).
Assume the setting of Theorem $1.$ Let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})\in K_{\mathsf{T}\times G_{\mathsf{v}}}(\mathrm{pt})$. Then
$V^{(\tau)}_{p}(z_{1},z_{2})=\sum_{F}\chi\left(F,\frac{(\operatorname{ev}_{0}^{*}(\tau_{1})\otimes\hat{\mathcal{O}}_{\operatorname{vir}})|_{F}}{\bigwedge(N_{\operatorname{vir}}|_{F}^{\vee})}\right)z_{1}^{\deg F}\iota^{*}V^{(\tau_{2}),\sigma_{F}}_{p^{(2)}}(z_{2})$
where the sum runs over $\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}$-fixed components $F$ of $\mathsf{QM}_{p^{(1)}}$
We prove this theorem by using equivariant localization, relating torus fixed quasimaps to $\mathcal{M}$ with torus fixed quasimaps to $\mathcal{M}^{(1)}$ and (twisted) quasimaps to $\mathcal{M}^{(2)}$. Theorem 1.2 can be thought of as a “branching rule” for vertex functions, describing a vertex function in terms of vertex functions for smaller quivers.
We explore several corollaries in Section 3.4 in which Theorem 3.2 can be made more transparent. In the special case when $\mathcal{M}^{(2)}$ is zero-dimensional, $V^{(1)}_{\mathcal{M}^{(2)}}$ will not depend on the so-called framing parameters. It follows that the twisted vertex is equal to the ordinary vertex up to a monomial in $z_{2}$, and we obtain the factorization
$V^{(\tau_{1}\otimes 1)}_{p}(z_{1},z_{2})=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})V^{(1)}_{p^{(2)}}(z_{2})$
where $z_{1}^{\prime}$ stands for a shift of $z_{1}$ by some powers of $z_{2}$, see Corollary 3.5.
1.4 Connection with quantum Hikita conjecture
As another special case, we can take the $q=\hbar$ specialization of Theorem 1.2, where $\hbar$ is the weight of the symplectic form of $\mathcal{M}$ and $q$ is the equivariant parameter for the action of $\mathbb{C}^{\times}$ on the source $\mathbb{P}^{1}$ of quasimaps. For certain descendants, this again forces each twisted vertex in Theorem 1.2 to be equal to the untwisted vertex up to a monomial in $z_{2}$. If we are lucky, this monomial can be absorbed into $z_{1}$, leading to another factorization
$V^{(\tau_{1}\otimes\tau_{2})}_{p}(z_{1},z_{2})\big|_{q=\hbar}=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})\big|_{q=\hbar}V^{(\tau_{2})}_{p^{(2)}}(z_{2})\big|_{q=\hbar},$ (2)
see Corollary 3.7 for the precise statement.
This is related to the quantum Hikita conjecture proposed in [22], which predicts an isomorphism between the $q=\hbar$ specialized quantum $D$-module for the Higgs branch $\mathcal{M}$ and the $D$-module of graded traces for the Coulomb branch.
A proof of this conjecture was given for ADE quiver varieties with minuscule framings in [15]. The technical heart of the proof is to identify the $q=\hbar$ specialized descendant vertex functions with the graded traces of elements of the quantized Coulomb branch–in this case a truncated shifted Yangian–over a Verma module. Thus our results here provide many cases of the Higgs side of the computation needed to prove the quantum Hikita for more general quiver varieties. In many examples, see Section 4.1, vertex functions can be computed in terms of reverse plane partitions over some poset. Theorem 1.2 shows that this is true for $V^{(\tau)}_{p}$ if it is true for $V^{(\tau_{1})}_{p^{(1)}}$ and $V^{(\tau_{2})}_{p^{(2)}}$.
Furthermore, passing (2), or more generally Theorem 1.2, through 3d mirror symmetry gives a conjectural “branching rule” for graded traces. To the best of the authors’ knowledge, this formula, and its expression in terms of reverse plane partitions in special cases, is a new expectation.
1.5 Conjecture for 0-dimensional vertex functions
Theorem 1.2 and our study of slant sums was motivated by 3d mirror symmetry. Let $\mathsf{A}\subset\mathsf{T}$ be the subtorus preserving the symplectic form and let $p\in\mathcal{M}^{\mathsf{A}}$. As proposed in Conjecture 1 in [16], the limit in the equivariant parameters $\mathsf{A}$ of the vertex $V_{p}$ should factorize to a product of $q$-binomials, one for each repelling weight of $T_{p^{!}}\mathcal{M}^{!}$, the tangent space of the mirror dual variety $\mathcal{M}^{!}$ at the dual fixed point $p^{!}$ [^2] It is also expected, and assumed here, that there is a bijection between fixed points on $\mathcal{M}$ and $\mathcal{M}^{!}$ when all fixed points are isolated.. This is a degeneration of the full 3d mirror symmetry of vertex functions proven in type $A$ in [5]. Conveniently, it can be checked even when the definition of vertex function of $\mathcal{M}^{!}$ is unknown, which is currently the case except in type $A$.
In particular, if $\mathcal{M}$ is a single point, the vertex function $V_{\mathcal{M}}$ is already independent of the equivariant parameters of $\mathsf{A}$ and is thus expected to factorize to a product of $q$-binomials. Here we make two conjectures refining the conjecture of [16], providing an explicit formula for the factorization and, equivalently, for the tangent space of $\mathcal{M}^{!}$.
To state it, we fix a quiver $Q$ without loops and a zero-dimensional quiver variety $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ where the stabilty parameter is $\theta=(1,1,\ldots,1)$. As usual, we define weights
$\lambda:=\sum_{i\in Q_{0}}\mathsf{w}_{i}\varpi_{i},\quad\mu:=\lambda-\sum_{i\in Q_{0}}\mathsf{v}_{i}\alpha_{i}$ (3)
of the Kac-Moody Lie algebra associated to the quiver $Q$. Here, $\varpi_{i}$ and $\alpha_{i}$ denote the fundamental weights and simple roots, respectively. Let $\Phi^{\pm,\text{re}}$ be the set of real positive or negative roots. Let $\Phi^{-,\text{re}}_{\mu}=\{\alpha\in\Phi^{-,\text{re}}\,\mid\,(\alpha,\mu)>0\}$ and $\Phi^{+,\text{re}}_{\mu}=\{\alpha\in\Phi^{+,\text{re}}\,\mid\,(\alpha,\mu)<0\}$. One can show, see Lemma 6.4, that $\dim\mathcal{M}=0\iff\mu\in W\lambda$.
The vertex function depends on a choice of polarization, and we choose the canonical polarization [^3] The is determined by the choice of orientation of the edges in the underlying graph of the quiver. Equivalently, we choose the polarization induced by the zero section of the cotangent bundle appearing in the symplectic reduction defining $\mathcal{M}$.. Then we define
$\mathsf{V}_{\mathcal{M}}(z)=V_{\mathcal{M}}(z)_{z=z(-\hbar^{-1/2})^{\det T^{1/2}}}$ (4)
i.e. the shift is
$z_{i}=z_{i}(-\hbar^{-1/2})^{a_{i}},\quad a_{i}=\sum_{\begin{subarray}{c}e\\ h(e)=i\end{subarray}}\mathsf{v}_{t(e)}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}+w_{i}$
Let $A_{Q}=(a_{i,j})_{i,j\in Q_{0}}$ be the adjacency matrix of the quiver, i.e., $a_{i,j}$ is the number of arrows from vertex $i$ to vertex $j$. We identify the Kähler parameters with formal exponentials of simple roots via $e^{\alpha_{i}}=(q/\hbar)^{(\mathsf{v}-A_{Q}\cdot\mathsf{v})_{i}}z_{i}$. More explicitly,
$e^{\alpha_{i}}=z_{i}(q/\hbar)^{b_{i}},\quad b_{i}=\mathsf{v}_{i}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}$
Let $\Phi(x):=\prod_{i=0}^{\infty}(1-xq^{i})$.
Conjecture 1.3.
The vertex function of $\mathcal{M}$ factorizes
$\mathsf{V}_{\mathcal{M}}(z)=\prod_{\begin{subarray}{c}\alpha\in\Phi^{+,\text{re}}_{\mu}\\ \end{subarray}}\prod_{i=1}^{-(\alpha,\mu)}\frac{\Phi\left(\hbar\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}{\Phi\left(\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}$
This conjecture has been proven for type $A$ quiver varieties with one framing by Smirnov and the first author in [17]. It was also proven by Jang and the first author for type $D$ quiver varieties with a single minuscule framing in [14]. Finally, for cotangent bundles of Grassmannians, it follows from [13]. One of our motivations in studying slant sums of quiver varieties was to develop tools to prove this conjecture more generally.
Equivalently, under the identification $e^{\alpha_{i}}=z_{i}(q/\hbar)^{-b_{i}}$, we have
$\mathsf{V}_{\mathcal{M}}(z^{-1})=\prod_{\begin{subarray}{c}\alpha\in\Phi^{-,\text{re}}_{\mu}\\ \end{subarray}}\prod_{i=1}^{\langle\alpha,\mu\rangle}\frac{\Phi\left(\hbar\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}{\Phi\left(\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}$ (5)
In this formula, we now view $\mu$ as a coweight using the same formulas as (3) but with fundamental coweights and simple coroots replacing fundamental weights and simple roots. We now write the pairing as $\langle\cdot,\cdot\rangle$ to reflect this. As discussed in [5], see also [1], this normalization of the vertex is the most natural for 3d mirror symmetry.
In §4.1, and specifically in Proposition 4.2, we demonstrate how Theorem 1.2 can be used to inductively prove instances of Conjecture 1.3.
Remark 1.4.
While this paper was in the final stages of preparation, we learned about a work in-progress by H. Nakajima, A. Okounkov, and Z. Zhou which contains a proof of Conjecture 1.3 for ADE quiver varieties (with no assumption about minuscule framings) using the universal deformation of quiver varieties [43]. The slant sum of two Dynkin quivers is almost never a Dynkin quiver; so the main interest of the techniques developed here lie outside ADE type.
1.6 Duals to zero-dimensional quiver varieties
In the same setting as the previous section, the 3d mirror dual to $\mathcal{M}$ is $\mathcal{M}^{!}=\overline{\mathcal{W}}^{\lambda}_{\mu}$, the corresponding affine Grassmannian slice. Combining (5) with Conjecture 1 from [16], and recalling that mirror symmetry identifies $\hbar^{!}=q/\hbar$, we obtain the following:
Conjecture 1.5.
The character of the tangent space of $\mathcal{M}^{!}$ at its unique torus fixed point $p^{!}$ is
$\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\sum_{i=1}^{\langle\alpha,\mu\rangle}(\hbar^{!})^{1-i}e^{\alpha}+(\hbar^{!})^{i}e^{-\alpha},$
We show in Lemma 6.9 that $\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\langle\alpha,\mu\rangle=\text{ht}(\lambda-\mu)=\sum_{i\in Q_{0}}\mathsf{v}_{i}$. The corresponding Coulomb branch $\mathcal{M}^{!}$ is birational to $T^{*}T_{\mathsf{v}}/W$, see [31], where $T_{\mathsf{v}}$ is a maximal torus of the gauge group $\prod_{i\in Q_{0}}GL(\mathsf{v}_{i})$ and $W$ is the Weyl group. Thus $\dim\mathcal{M}^{!}=2|\mathsf{v}|$, in agreement with Conjecture 1.5. This numerical coincidence provides some numerical explanation for the appearance of only real roots. It would be desirable to have a more conceptual explanation.
In Section 6, we will prove the following.
Proposition 1.6 (Proposition 6.7).
Conjecture 1.5 holds in ADE type.
1.7 Partial resolutions of Coulomb branches
We expect that Conjecture 1.5 also provides a formula for the tangent space to an arbitrary partial resolution of a Coulomb branch at a nonsingular fixed point. In Section 5.4, we recall some facts about partial resolutions of Coulomb branches. We conjecture that isolated fixed points on the Higgs branch correspond to nonsingular torus fixed points on a partial resolution of the Coulomb branch. Furthermore, this correspondence works in such a way that the character of the tangent space at an arbitrary nonsingular fixed point on a partial resolution can be extracted from Conjecture 1.5, see Corollary 5.4.
1.8 Modules over quantized Coulomb branches
Assume that Conjecture 1.5 holds. We observe in Section 5.4 that the category $\mathcal{O}$ for a quantization of $\mathcal{M}^{!}$ contains a unique irreducible object and conjecture that its character can be read off from the explicit formula for the tangent space to the Coulomb branch at a fixed point (see Conjecture 5.7). More generally, we believe that Conjecture 1.5 should provide a formula for normalized characters of a natural family of modules over Coulomb branches labeled by nonsingular fixed points of their partial resolutions, see Remark 5.9 for a speculation about this.
1.9 Slant sums of Coulomb branches
In Section 7, we initiate the study of the slant sum construction from the Coulomb branch perspective. We prove that for $\mathsf{w}^{(2)}$ such that $\mathsf{w}^{(2)}_{\star_{2}}=1$ and all other $\mathsf{w}_{i}^{(2)}=0$, the slant sum on the Coulomb branch side corresponds to taking products (see Proposition 7.2). Note that the identification of a Coulomb branch $\mathcal{M}_{C}$ for a quiver $Q$ with the product $\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)}$ of Coulomb branches for $Q^{(1)},Q^{(2)}$ is compatible with natural structures that we have (integrable system, torus action) but only after a certain twist.
We also mention a partial description of the Coulomb branch for the slant sum for arbitrary $\mathsf{w}^{(2)}$ such that $\mathsf{w}^{(2)}_{\star_{2}}=1$ (see Proposition 7.7 that is an immediate corollary of [8, Proposition 3.18]). For $\mathsf{w}^{(2)}_{\star_{2}}>1$, we do not know how exactly $\mathcal{M}_{C}$, $\mathcal{M}_{C}^{(1)}$, $\mathcal{M}_{C}^{(2)}$ are related.
On the other hand, recall that the $\hbar=q$ specialized vertex functions restricted to fixed points are expected to equal graded traces of Verma modules over quantized Coulomb branches. So, our Higgs side computation suggests not a relation between (quantized) Coulomb branches for $Q,Q^{(1)}$, and $Q^{(2)}$ but between modules over some twisted versions of them.
Namely, one can consider the category of so-called Gelfand-Tsetlin modules over the quantized Coulomb branch $\mathcal{A}_{Q}$, see [42]. Let ${\mathsf{A}}_{\mathsf{v}}\subset G_{\mathsf{v}}$ be a maximal torus. It follows from [42] (see also [33, Appendix B] and [41] for the geometric treatment) that “blocks” of categories of GT-modules over $\mathcal{A}_{Q}$ are equivalent to categories of (topologically nilpotent) modules over the convolution algebra $\widehat{H}_{*}^{Z_{G}(\sigma_{\mathsf{v}})\times\mathsf{A}}(\mathcal{T}_{Q}^{(\sigma(t),t)}\times_{{\bf{N}}_{\mathcal{K}}^{(\sigma(t),t)}}\mathcal{T}_{Q}^{(\sigma(t),t)})$. Such a block depends on a choice of a cocharacter $\sigma\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}_{\mathsf{v}}\times\mathsf{A}$. Here $\mathcal{T}_{Q}=G_{\mathcal{K}}\times^{G_{\mathcal{O}}}{\bf{N}}_{\mathcal{O}}$. Fixed points $\mathcal{T}_{Q}^{(\sigma(t),t)}$ can be explicitly computed and turn out to be equal to the disjoint unions of products of the corresponding fixed points on $\mathcal{T}_{Q^{(1)}}$ as well as fixed points on some twisted version of $\mathcal{T}_{Q^{(2)}}$, this is completely parallel to Theorem 3.1, where we provide a similar description of torus fixed based quasimaps to the resolved Higgs branch for $Q$. As a corollary of the above, one should be able to recover a branching formula for a graded trace of a GT-module over $\mathcal{A}_{Q}$ reflecting our Theorem 1.2 on the Coulomb side. We will return to this in the second version of the paper.
1.10 Branching for non-stationary Ruijsenaars function
In addition to providing a tool with which to study the vertex functions of zero-dimensional quiver varieties, Theorem 1.2 also subsumes and was motivated by an additional formula present in the literature, which we discuss in detail in Section 4.2. Let $X_{n}$ be the cotangent bundle to the variety of complete flags in $\mathbb{C}^{n}$. Viewed as a quiver variety, it decomposes as the slant sum of a cotangent bundle of a Grassmannian and a smaller flag variety: $X_{n}=T^{*}Gr(n-1,n)\#X_{n-1}$.
The vertex function of $X_{n}$ is the so-called “nonstationary Ruijsenaars function” studied, for example, in [9, 26, 35, 40]. It has the remarkable property of being solutions to the Macdonald difference equations in both the equivariant parameters for $\mathsf{A}$ and the Kähler variables.
Then formula (3.11) of [35] (see also (5.21)) is exactly Theorem 1.2 in this case. Furthermore, Theorem 1.2 also gives a similar branching formula for the cotangent bundles of partial flag varieties.
1.11 Outline
In Section 2, we establish notation and review the facts we need about quiver varieties. We also clearly state the exact hypotheses needed in Theorem 1.1 and prove this theorem.
In Section 3, we review vertex functions, prove Theorem 1.2, and discuss a number of factorization corollaries.
In Section 4, we consider some examples which show how Theorem 1.2 relates to Conjecture 1.3. We also explicitly spell out the connection between our results the branching of the non-stationary Ruijsenaars function of [35].
In Section 5, we recall some facts about Higgs and Coulomb branches and their mirror symmetry. We make some conjectures needed for Conjecture 1.5 to make sense. We also make some conjectures which would reduce the character of tangent spaces at fixed points for an arbitrary (partial resolution of a) Coulomb branch to the case of Conjecture 1.5.
In Section 6, we prove all of the conjectures in ADE types.
In Section 7, we investigate slant sums from the perspective of Coulomb branches.
1.12 Acknowledgements
We are grateful to Jiwu Jang, Shrawan Kumar, and Zijun Zhou for helpful conversations regarding this project. H. Dinkins was supported by NSF grant DMS-2303286 at MIT and the NSF RTG grant Algebraic Geometry and Representation Theory at Northeastern University DMS–1645877. R. Lance was supported by NSF grant DMS-2200867. Vasily Krylov was supported by the Simons Foundation Award 888988 as part of the Simons Collaboration on Global Categorical Symmetries.
2 Slant sums of Higgs branches
We briefly recall definitions, theorems, and fix conventions regarding quiver varieties. A more detailed introduction can be found in [32].
A quiver, $Q$, is a directed graph. It comes equipped with a finite vertex set, $Q_{0}$ and a set of directed edges $Q_{1}$. There are maps $h,t:Q_{1}\to Q_{0}$ sending an arrow to its head and tail. A representation of the quiver $Q=(Q_{0},Q_{1})$ of dimension $\mathsf{v}\in\mathbb{N}^{Q_{0}}$ is a collection of complex vector spaces $V_{i}$ with $\dim V_{i}=\mathsf{v}_{i}$ for each $i\in Q_{0}$ and linear maps $x_{e}:V_{t(e)}\to V_{h(e)}$ for each arrow $e\in Q_{1}$.
Choose $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$ and fix a pair of $Q_{0}$-graded vector spaces $V=\bigoplus_{i\in Q_{0}}V_{i},W=\bigoplus_{i\in Q_{0}}W_{i}$ such that $\mathsf{v}_{i}=\dim V_{i}$ and $\mathsf{w}_{i}=\dim W_{i}$. Let
${\bf{M}}:={\bf{M}}(\mathsf{v},\mathsf{w}):=\Bigg(\bigoplus_{e\in Q_{1}}\operatorname{Hom}(V_{t(e)},V_{h(e)})\Bigg)$ $\oplus\Bigg(\bigoplus_{e\in Q_{1}}\operatorname{Hom}(V_{h(e)},V_{t(e)})\Bigg)$
$\oplus\Bigg(\bigoplus_{i\in Q_{0}}\operatorname{Hom}(W_{i},V_{i})\Bigg)$ $\oplus\Bigg(\bigoplus_{i\in Q_{0}}\operatorname{Hom}(V_{i},W_{i})\Bigg)$
which, by the trace pairing, may be identified with $T^{*}{\bf{N}}$, where
${\bf{N}}:={\bf{N}}(\mathsf{v},\mathsf{w}):=\bigoplus_{e\in Q_{1}}\operatorname{Hom}(V_{t(e)},V_{h(e)})\oplus\bigoplus_{i\in Q_{0}}\operatorname{Hom}(V_{i},W_{i})$ (6)
We denote an element of $\bf{M}$} as $(X,Y,I,J)$, where
$X=\bigoplus_{e\in Q_{1}}X_{e},\quad Y=\bigoplus_{e\in Q_{1}}Y_{e},\quad I=\bigoplus_{i\in Q_{0}}I_{i},\quad J=\bigoplus_{i\in Q_{0}}J_{i}$
As the cotangent bundle to a smooth manifold, $\bf{M}$} comes with a canonical symplectic form.
$\bf{M}$} has an action of $G_{\mathsf{v}}:=\prod_{i\in Q_{0}}GL(V_{i})$, sometimes referred to as the gauge group, induced via basis change at each vertex:
$g.(X,Y,I,J)=(gXg^{-1},gYg^{-1},gI,Jg^{-1})$
We denote $\text{Lie}(G_{\mathsf{v}}):=\mathfrak{g}_{\mathsf{v}}$.
The action of $G_{\mathsf{v}}$ on $\bf{M}$} is Hamiltonian, and the moment map
$\mu:{\bf{M}}\to\mathfrak{g}_{\mathsf{v}}$
is given by
$(X,Y,I,J)\mapsto\bigoplus_{i\in Q_{0}}\left(\sum_{\begin{subarray}{c}e\in Q_{1}\\ h(e)=i\end{subarray}}X_{e}Y_{e}-\sum_{\begin{subarray}{c}e\in Q_{1}\\ t(e)=i\end{subarray}}Y_{e}X_{e}+I_{i}J_{i}\right)$
where we have identified $\mathfrak{g}_{\mathsf{v}}\cong\mathfrak{g}_{\mathsf{v}}^{*}$ via the trace pairing.
A choice of stability parameter $\theta\in\mathbb{Z}^{Q_{0}}$ induces a character $\chi_{\theta}$ of $G_{\mathsf{v}}$ by $\chi_{\theta}(g)=\prod_{i\in Q_{0}}(\det g_{i})^{\theta_{i}}$.
Definition 2.1.
The Nakajima quiver variety associated to the data $Q$, $V_{i}$, $W_{i}$, and $\theta$ is the algebraic symplectic reduction
$\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})={\bf{M}}/\!\!/\!\!/_{\chi_{\theta}}G_{\mathsf{v}}=\mu^{-1}(0)/\!\!/_{\chi_{\theta}}G_{\mathsf{v}}$
We may omit $Q,\theta$ when they are understood. By the general theory of GIT, $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is a quasiprojective variety and admits a projective morphism
$\pi:\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\to\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w})$
to the affine categorical quotient $\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w})=\mu^{-1}(0)/\!\!/G_{\mathsf{v}}$.
Proposition 2.2 ([30]).
Fix $Q$, $\mathsf{v}$, and $\mathsf{w}$. There is a finite set of hyperplanes in $\mathbb{Z}^{Q_{0}}\otimes_{\mathbb{Z}}\mathbb{R}$ such that if $\theta$ is not on a hyperplane, then $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is nonsingular, symplectic, and connected.
We will call stability parameters as in the previous proposition as “generic”. For this paper it suffices to know that $\theta=\pm(1,\dots,1)$ is always generic.
Furthermore, GIT provides a notion of $\theta$ (semi)stable points. It is known that for generic $\theta$, stability and semistability are equivalent. Thus the closed points of $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ are in bijection with orbits $\mu^{-1}(0)^{\theta-s}/G_{\mathsf{v}}$. The superscript here refers to the intersection of $\mu^{-1}(0)$ with the locus of $\theta$-stable points. The following proposition gives a criteria for stability.
Proposition 2.3 ([32]).
Fix $\theta\in\mathbb{Z}^{Q_{0}}$. A representation $V=(X,Y,I,J)\in{\bf{M}}$ is $\theta$-semistable if for any proper nonzero subrepresentation, $V^{\prime}$, we have
$(V^{\prime}\subset\ker J)$ $\Rightarrow\theta\cdot\dim V^{\prime}\geq 0$
$(V^{\prime}\supset\operatorname{im}I)$ $\Rightarrow\theta\cdot\dim V^{\prime}\geq\theta\cdot\dim V$
and $\theta$-stable if both implications hold with strict inequalities.
It follows from Proposition 2.3 that all stability conditions $\theta$ such that $\theta_{i}>0$ for all $i\in Q_{0}$ are equivalent to each other. We will denote such a stability condition by $\theta>0$. Similarly, we have stability conditions such that $\theta<0$.
Under some assumptions, $\pi:\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\to\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w})$ is a symplectic resolution of singularities of its image.
When the data $Q$, $\mathsf{v}$, $\mathsf{w}$, and $\theta$ is fixed, we will often save on notation and denote $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$.
2.1 Torus action and fixed points
Fix $Q$, $\mathsf{v}$, $\mathsf{w}$, and $\theta$. The vector space $\bf{M}$} has actions of the framing torus $\mathsf{A}:=(\mathbb{C}^{\times})^{|\mathsf{w}|}$. There is another action of a rank one torus, denoted $\mathbb{C}^{\times}_{\hbar}$ by scaling the cotangent fibers with weight $1$. In formulas,
$(a,\hbar).(X,Y,I,J)=(X,\hbar^{-1}Y,Ia^{-1},\hbar^{-1}aJ),\quad(a,\hbar)\in\mathsf{A}\times\mathbb{C}^{\times}_{\hbar}$
Both actions descend to $\mathcal{M}$, and we denote $\mathsf{T}:=\mathsf{A}\times\mathbb{C}^{\times}_{\hbar}$. We will be interested in the $\mathsf{T}$-fixed locus of $\mathcal{M}$.
The vector spaces $V_{i}$ descend to tautological bundles $\mathcal{V}_{i}$ over $\mathcal{M}$, defined by
$\mathcal{V}_{i}=(\mu^{-1}(0)\times V_{i})/G_{\mathsf{v}}$
where $g\in G_{\mathsf{v}}$ acts on $(p,v)\in\mu^{-1}(0)\times V_{i}$ by $g\cdot(p,v)=(g^{-1}p,g_{i}^{-1}v)$. Similarly, there are topologically trivial tautological bundles $\mathcal{W}_{i}$ induced by the framing spaces. Both types of bundles are $\mathsf{T}$-equivariant.
When $p\in\mathcal{M}^{\mathsf{T}}$, the vector space $\mathcal{V}_{i}|_{p}$ is a $\mathsf{T}$-module, and thus may be decomposed into $\mathsf{T}$ weight spaces.
Let $p\in\mathcal{M}^{\mathsf{T}}$ and let $[(X,Y,I,J)]$ be a representative of $p$. Since $p$ is $\mathsf{T}$-fixed, there is a homomorphism $\rho:\mathsf{T}\to G_{\mathsf{v}}$ such that
$(a,\hbar)\cdot(X,Y,I,J)=\rho(a,\hbar)\cdot(X,Y,I,J)$
Written explicitly,
$(X,\hbar^{-1}Y,Ia^{-1},\hbar^{-1}aJ)=(gXg^{-1},gYg^{-1},gI,Jg^{-1}),\quad g=\rho(a,\hbar)$
So $\mathsf{T}$ acts on each $V_{i}$ space via $\rho$. Let $a_{i,j}$ for $i\in Q_{0}$ and $1\leq j\leq\mathsf{w}_{i}$ be the $\mathsf{A}$-weights of the framing spaces $W_{i}$. The decomposition of $V_{i}$ into $\mathsf{T}$-weight spaces takes the form
$V_{i}=\bigoplus_{j\in Q_{0}}\bigoplus_{k=1}^{\mathsf{w}_{j}}\bigoplus_{l\in\mathbb{Z}}V_{i}(a_{j,k}^{\pm}\hbar^{l})$
where $V_{i}(a_{j,k}\hbar^{l})$ denotes the (possibly empty) space of weight $a_{j,k}\hbar^{l}$ of $V_{i}$.
From the equation $gI=Ia^{-1}$, it follows that $I(W_{i}(a_{i,j}))\subset V(a_{ij}^{-1})$. Similarly, the equation $Jg^{-1}=\hbar^{-1}aJ$ implies $J(V_{i}(w))\subset W_{i}(\hbar w^{-1})$. Furthermore, for any $e\in Q_{1}$, we have $X(V_{t(e)}(w))\subset V_{h(e)}(w)$ and $Y(V_{h(e)}(w))\subset V_{t(e)}(\hbar^{-1}w)$.
Note also that the $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{p}$ are the inverse of the weights of $V_{i}$ under the $\mathsf{T}$-action by $\rho$ for the fixed point $p$.
We will need the following lemmas characterizing representatives of equivalence classes of $\mathsf{T}$-fixed stable representations. They utilize the computation and notation from above.
Lemma 2.4.
Consider a quiver variety $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$, and let $p\in\mathcal{M}^{\mathsf{T}}$. Let $V$ be a representative of $p$. Then
  1. 1.
    If $\theta<0$, then $V_{i}(a_{j,k}^{b}\hbar^{c})=0$ unless b=1b=-1 and $c\geq 1$ for all $i\in Q_{0}$.
  2. 2.
    If $\theta>0$, then $V_{i}(a_{j,k}^{b}\hbar^{c})=0$ unless b=1b=-1 and $c\leq 0$ for all $i\in Q_{0}$..
Proof.
By Proposition 2.3, if $\theta>0$, then for a stable point $(X,Y,I,J)$, the image of the $I$ generates each $V_{i}$ under the action of $X$ and $Y$. Part 2 follows by the above discussion of how the quiver maps affect weight spaces. Part 1 is similar.
Lemma 2.5.
Consider a quiver variety $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$, and let $p\in\mathcal{M}^{\mathsf{T}}$. Let $(X,Y,I,J)$ be a representative of $p$.
  1. 1.
    $J_{i}I_{i}=0$ for all $i\in Q_{0}$.
  2. 2.
    If $\theta<0$, then $I_{i}=0$ for all $i\in Q_{0}$.
  3. 3.
    If $\theta>0$, then $J_{i}=0$ for all $i\in Q_{0}$.
Proof.
Since $p$ is $\mathsf{T}$-fixed, $JI(W_{i}(a_{i,j}))\subset W_{i}(a_{i,j}\hbar)$. Since framing spaces have trivial $\mathbb{C}^{\times}_{\hbar}$-weight, $JI=0$.
For statements 2 and 3, we utilize Lemma 2.4.
Since $p$ is $\mathsf{T}$-fixed, $I_{i}(W_{i}(a_{i,j}))\subset V_{i}(a_{i,j}^{-1})$. If $\theta<0$, the latter weight space is 0 by Lemma 2.4. This proves statement 2.
Similarly, $J_{i}(V_{i}(a_{j,k}^{-1}\hbar^{c}))\subset W_{i}(a_{j,k}\hbar^{1-c})$. If $\theta>0$, the former weight space is 0 unless $c\leq 0$, in which case the latter is 0. This proves statement 3. ∎
2.2 Slant sum of quivers
Let $Q^{(1)}$ and $Q^{(2)}$ be quivers and let $\star_{r}\in Q^{(r)}_{0}$. Let $Q$ be the quiver $Q=(Q_{0},Q_{1})$ such that $Q_{0}=Q^{(1)}_{0}\sqcup Q^{(2)}_{0}$ and $Q_{1}=Q^{(1)}_{1}\sqcup Q^{(2)}_{1}\sqcup\{\star_{1}\to\star_{2}\}$. We call $Q$ the slant sum of $Q^{(1)}$ and $Q^{(2)}$ over the vertices $\star_{1}$ and $\star_{2}$ and will denote this by
$Q=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$
or sometimes just $Q=Q^{(1)}\#Q^{(2)}$ if the choices of $\star_{1}$ and $\star_{2}$ are clear from context. In words, $Q$ is given by taking the disjoint union of $Q^{(1)}$ and $Q^{(2)}$ and adding an arrow from $\star_{1}$ to $\star_{2}$.
Let $\mathsf{v}^{(r)},\mathsf{w}^{(r)}\in\mathbb{Z}_{\geq 0}^{Q^{(r)}_{0}}$. We call $\star_{1}$ and $\star_{2}$ compatible if $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}$.
Let $\mathsf{v},\mathsf{w}\in\mathbb{Z}_{\geq}^{Q_{0}}$ be defined by
$\mathsf{v}_{i}=\begin{cases}\mathsf{v}^{(1)}_{i}&\text{if $i\in Q^{(1)}_{0}$}\\ \mathsf{v}^{(2)}_{i}&\text{if $i\in Q^{(2)}_{0}$}\end{cases}$
and
$\mathsf{w}_{i}=\begin{cases}\mathsf{w}^{(r)}_{i}&\text{if $i\in Q^{(r)}_{0}\setminus\{\star_{2}\}$, $r\in\{1,2\}$}\\ 0&\text{if $i=\star_{2}$}\end{cases}$
We will denote this by
$\mathsf{v}=\mathsf{v}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathsf{v}^{(2)},\quad\mathsf{w}=\mathsf{w}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathsf{w}^{(2)}$
or sometimes just $\mathsf{v}^{(1)}\#\mathsf{v}^{(2)}$ (similarly for $\mathsf{w}$).
Intuitively, we have identified the gauge vertex at $\star_{1}$ with the framed vertex at $\star_{2}$ and chosen to view the result as a gauge vertex. All together, we call $Q,\mathsf{v},\mathsf{w}$ the slant sum of the quiver gauge theories associated to $Q^{(1)},\mathsf{v}^{(1)},\mathsf{w}^{(1)}$ and $Q^{(2)},\mathsf{v}^{(2)},\mathsf{w}^{(2)}$.
Finally, given stability vectors $\theta^{(r)}\in\mathbb{Z}^{Q^{(r)}_{0}}$, we define $\theta\in\mathbb{Z}^{Q_{0}}$ by
$\theta_{i}=\theta^{(r)}_{i},\quad\text{ where $i\in Q^{(r)}_{0}$}$
and denote it as $\theta^{(1)}\#\theta^{(2)}$.
Definition 2.6.
Fix the data of $Q^{(r)}$, $\mathsf{v}^{(r)}$, $\mathsf{w}^{(r)}$, and $\theta^{(r)}$ as above for $r\in\{1,2\}$. Let $\star_{1}$ and $\star_{2}$ be compatible vertices. Define $Q$, $\mathsf{v}$, $\mathsf{w}$, and $\theta$ as above. The slant sum of $\mathcal{M}_{Q^{(1)},\theta^{(1)}}(\mathsf{v}^{(1)},\mathsf{w}^{(1)})$ and $\mathcal{M}_{Q^{(2)},\theta^{(2)}}(\mathsf{v}^{(1)},\mathsf{w}^{(2)})$ over $\star_{1}$ and $\star_{2}$ is the quiver variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$.
We will at times omit pieces of the notation and denote this by $\mathcal{M}^{(1)}\#\mathcal{M}^{(2)}$.
2.3 Relationship between slant sums
Fix $\mathcal{M}^{(r)}:=\mathcal{M}_{Q^{(r)},\theta^{(r)}}(\mathsf{v}^{(r)},\mathsf{w}^{(r)})$ for $r\in\{1,2\}$ as above. Let $\star_{1}$ and $\star_{2}$ be compatible vertices. Let $\mathcal{M}:=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$. There is a diagram relating these quiver varieties.
We fix the vector spaces $V^{(r)}_{i}$ and $W^{(r)}_{i}$ used in the construction of $\mathcal{M}^{(r)}$. Then we use
$V_{i}:=V^{(r)}_{i}\text{ where $i\in Q^{(r)}_{0}$}$
and
$W_{i}:=\begin{cases}W^{(r)}_{i}&\text{if $i\in Q^{(r)}_{0}\setminus\{\star_{2}\}$, $r\in\{1,2\}$}\\ 0&\text{if $i=\star_{2}$}\end{cases}$
as the vector spaces in the construction of $\mathcal{M}$.
Consider the affine variety
${\bf{M}}(\mathsf{v}^{(1)},\mathsf{w}^{(1)})\times\operatorname{Hom}(V^{(1)}_{\star_{1}},W^{(2)}_{\star_{2}})_{\text{iso}}\times{\bf{M}}(\mathsf{v}^{(2)},\mathsf{w}^{(2)})$
where the subscript iso denotes the subvariety consisting of linear isomorphisms. Let $\bf{Y}$ be the subvariety satisfying the moment map equations and stability for both $Q^{(1)}$ and $Q^{(2)}$, i.e.
${\bf{Y}}=\mu^{-1}_{1}(0)^{s}\times\operatorname{Hom}(V^{(1)}_{\star_{1}},W^{(2)}_{\star_{2}})_{\text{iso}}\times\mu^{-1}_{2}(0)^{s}$
The projection
${\bf{Y}}\to\mu^{-1}_{1}(0)^{s}\times\mu^{-1}_{2}(0)^{s}$
is a $G=G_{\mathsf{v}^{(1)}}\times G_{\mathsf{v}^{(2)}}$-equivariant map which thus descends to the quotients
${\bf{Y}}/G\to\mathcal{M}^{(1)}\times\mathcal{M}^{(2)}$
We denote points in $\bf{Y}$} by $(p^{(1)},\phi,p^{(2)})$, where $p^{(r)}=(X^{(r)},Y^{(r)},I^{(r)},J^{(r)})$ for $r\in\{1,2\}$ denotes the quiver representation as usual. We define a morphism
$F:{\bf{Y}}\to{\bf{M}}(\mathsf{v},\mathsf{w})$
by
$F(p^{(1)},\phi,p^{(2)})=(X,Y,I,J)$
where the components of the linear maps $(X,Y,I,J)$ are defined as follows. Set
$X_{e}=X^{(r)}_{e}\quad\text{where $e\in Q^{(r)}_{1}$}$
and similarly for $Y_{e}$. Set
$I_{i}=\begin{cases}I^{(r)}_{i}&\text{if $i\in Q^{(r)}_{0}\setminus\{\star_{2}\}$}\\ 0&\text{if $i=\star_{2}$}\end{cases}$
and similarly for $J_{i}$. Finally, for the new edge $e_{0}=\star_{1}\to\star_{2}$, we set $X_{e_{0}}=I^{(2)}_{\star_{2}}\circ\phi$ and $Y_{e_{0}}=\phi^{-1}\circ J^{(2)}_{\star_{2}}$.
Lemma 2.7.
The morphism $F$ is $G$-equivariant.
Proof.
The only things to check are that $g_{\star_{2}}X_{e_{0}}g_{\star_{1}}^{-1}=g_{\star_{2}}I^{(2)}_{\star_{2}}\phi^{\prime}$ and $g_{\star_{1}}Y_{e_{0}}g_{\star_{2}}^{-1}=\phi^{\prime-1}J^{(2)}_{\star_{2}}g_{\star_{2}}^{-1}$ where $\phi^{\prime}=\phi g_{\star_{1}}^{-1}$. These are immediate. ∎
Let ${\bf{Z}}=F^{-1}(\mu^{-1}(0)^{s})\subset{\bf{Y}}$.
By the definitions, we obtain a diagram
$\mu^{-1}_{1}(0)^{s}\times\mu^{-1}_{2}(0)^{s}$$\bf{Y}$$\bf{Z}$$\mu^{-1}(0)^{s}$$\mathcal{M}^{(1)}\times\mathcal{M}^{(2)}$${\bf{Y}}/G$${\bf{Z}}/G$$\mathcal{M}$} (7)
Later we will need the following lemma.
Lemma 2.8.
The following holds:
  1. 1.
    The closed subset of $\bf{Y}$} defined by $J^{(2)}_{\star_{2}}I^{(2)}_{\star_{2}}=0$ is equal to $F^{-1}(\mu^{-1}(0))$.
  2. 2.
    If $\theta^{(2)}<0$, then $\{I^{(2)}_{\star_{2}}=0\}\subset{\bf{Z}}$.
  3. 3.
    If $\theta^{(2)}>0$, then $\{J^{(2)}_{\star_{2}}=0\}\subset{\bf{Z}}$.
Proof.
Consider the first statement. The moment map equations for $\mathcal{M}$ differ from those of $\mathcal{M}^{(1)}$ or $\mathcal{M}^{(2)}$ only at the two vertices $\star_{1}$ and $\star_{2}$. The $\star_{1}$-component of the moment map for the $G$ action on ${\bf{M}}(\mathsf{v},\mathsf{w})$ applied to $F(p^{(1)},\phi,p^{(2)})$ is
$\sum_{\begin{subarray}{c}e\in Q_{1}\\ h(e)=\star_{1}\end{subarray}}X_{e}Y_{e}-\sum_{\begin{subarray}{c}e\in Q_{1}\\ t(e)=\star_{1}\end{subarray}}Y_{e}X_{e}+I_{\star_{1}}J_{\star_{1}}$
$=\sum_{\begin{subarray}{c}e\in Q^{(1)}_{1}\\ h(e)=\star_{1}\end{subarray}}X^{(1)}_{e}Y^{(1)}_{e}-\sum_{\begin{subarray}{c}e\in Q^{(1)}_{1}\\ t(e)=\star_{1}\end{subarray}}Y^{(1)}_{e}X^{(1)}_{e}+I^{(1)}_{\star_{1}}J^{(1)}_{\star_{1}}-Y_{e_{0}}X_{e_{0}}$
$=-Y_{e_{0}}X_{e_{0}}$
$=\phi^{-1}J^{(2)}_{\star_{2}}I^{(2)}_{\star_{2}}\phi$
Similarly, the $\star_{2}$-component of the moment map is
$\sum_{\begin{subarray}{c}e\in Q_{1}\\ h(e)=\star_{2}\end{subarray}}X_{e}Y_{e}-\sum_{\begin{subarray}{c}e\in Q_{1}\\ t(e)=\star_{2}\end{subarray}}Y_{e}X_{e}$
$=\sum_{\begin{subarray}{c}e\in Q^{(2)}_{1}\\ h(e)=\star_{2}\end{subarray}}X^{(2)}_{e}Y^{(2)}_{e}-\sum_{\begin{subarray}{c}e\in Q^{(2)}_{1}\\ t(e)=\star_{2}\end{subarray}}Y^{(2)}_{e}X^{(2)}_{e}+X_{e_{0}}Y_{e_{0}}$
$=\sum_{\begin{subarray}{c}e\in Q^{(2)}_{1}\\ h(e)=\star_{2}\end{subarray}}X^{(2)}_{e}Y^{(2)}_{e}-\sum_{\begin{subarray}{c}e\in Q^{(2)}_{1}\\ t(e)=\star_{2}\end{subarray}}Y^{(2)}_{e}X^{(2)}_{e}+I^{(2)}_{\star_{2}}J^{(2)}_{\star_{2}}$
which is exactly the $\star_{2}$-component of the moment map equation for the $G_{\mathsf{v}^{(2)}}$ action on ${\bf{M}}(\mathsf{v}^{(2)},\mathsf{w}^{(2)})$ and hence is 0. This proves the first statement.
For the second statement, recall the representation-theoretic characterization of stability from Proposition 2.3.
Let $S=(S_{i})_{i\in Q_{0}}$ be a subrepresentation of
$f(p^{(1)},\phi,p^{(2)})=(X,Y,I,J)$
such that $S_{i}\subseteq\ker J_{i}$ for all $i\in Q_{0}$.
There is a decomposition
$(S_{i})_{i\in Q_{0}}=(S^{(1)}_{i})_{i\in Q_{0}^{(1)}}\oplus(S_{i}^{(2)})_{i\in Q_{0}^{(2)}}.$
Denote $S^{(r)}=(S_{i}^{(r)})_{i\in Q_{0}^{(r)}}$ for $r=1,2$. It is automatic that $S^{(1)}$ is a subrepresentation of $(X^{(1)},Y^{(1)},I^{(1)},J^{(1)})$ for which $S^{(1)}_{i}\subseteq\ker J^{(1)}_{i}$ for all $i\in Q_{0}^{(1)}$. So stability for $p^{(1)}$ implies $\theta^{(1)}\cdot\dim S^{(1)}<0$. If $\theta^{(2)}=-(1,1,\ldots,1)$, then we also have $\theta^{(2)}\cdot\dim S^{(2)}\leq 0$. If $\theta^{(2)}=(1,1,\ldots,1)$ and $J^{(2)}_{\star_{2}}=0$, then $S^{(2)}$ is also a subrepresentation of $p^{(2)}$ for which $S^{(2)}_{i}\subseteq\ker J^{(2)}_{i}$ for all $i\in Q_{0}^{(2)}$; so stability for $p^{(2)}$ implies $\theta^{(2)}\cdot\dim S^{(2)}\leq 0$. In either case, we have
$\theta\cdot\dim S=\theta^{(1)}\cdot\dim S^{(1)}+\theta^{(2)}\cdot\dim S^{(2)}\leq 0$
Now let $S=(S_{i})_{i\in Q_{0}}$ be a subrepresentation of $(X,Y,I,J)$ such that $\operatorname{im}I_{i}\subseteq S_{i}$ for all $i\in Q_{0}$. As above, we obtain $S^{(1)}$ and $S^{(2)}$. It is automatic that $\operatorname{im}I^{(1)}_{i}\subseteq S^{(1)}_{i}$ for all $i\in Q^{(1)}_{i}$. So stability for $p^{(1)}$ implies that $\theta^{(1)}\cdot\dim S^{(1)}\leq\theta\cdot\dim V^{(1)}$. If $\theta^{(2)}=(1,1,\ldots,1)$, then $\theta^{(2)}\cdot\dim S^{(2)}\leq\theta^{(2)}\dim V^{(2)}$. If $\theta^{(2)}=-(1,1,\ldots,1)$ and $I_{\star_{2}}^{(2)}=0$, then $S^{(2)}$ is also a subrepresentation of $p^{(2)}$ such that $\operatorname{im}I^{(2}_{i}\subseteq S^{(2)}_{i}$ for all $i\in Q^{(2)}_{0}$; so stability for $p^{(2)}$ implies $\theta^{(2)}\cdot\dim S^{(2)}\leq\theta^{(2)}\cdot\dim V^{(2)}$. Either way
$\theta\cdot\dim S=\theta^{(1)}\cdot\dim S^{(1)}+\theta^{(2)}\cdot\dim S^{(2)}\leq\theta\cdot\dim V$
This concludes the proofs of parts 2 and 3.
2.4 Split fixed points
We next define the notion of split fixed points. This is a technical assumption that allows us to choose a (almost) canonical basis of a gauge vector space, allowing us in practice to view it as a framing space.
Definition 2.9.
Let $\mathcal{M}$ be a quiver variety. Let $p\in\mathcal{M}^{\mathsf{T}}$ be a fixed point. We say that $p$ is split over $i\in Q_{0}$ if all the $\mathsf{T}$-weight spaces of $\mathcal{V}_{i}|_{p}$ have dimension at most $1$.
Equivalently, by the discussion of Section 2.1, choosing a representative $\tilde{p}$ of the fixed point $p$, there is a homomorphism $\rho:\mathsf{T}\to G_{\mathsf{v}}$ such that $t\cdot\tilde{p}=\rho(t)\cdot\tilde{p}$ and the joint eigenspaces of the action of $\rho(\mathsf{T})$ on the vector space $V_{i}$ are all one dimensional.
Definition 2.10.
Let $p\in\mathcal{M}^{\mathsf{T}}$ be a split fixed point over $i\in Q_{0}$. A $(p,i)$-chamber of $\mathsf{T}$, denoted $\mathfrak{C}$, is a choice of total ordering of the $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{p}$.
Choosing a weight basis with respect to the ordering given by a $(p,i)$-chamber $\mathfrak{C}$, we obtain a canonical isomorphism $V_{i}\cong\mathbb{C}^{\mathsf{v}_{i}}$. Different choices of basis give isomorphisms that differ by a diagonal matrix in $GL(\mathbb{C}^{\mathsf{v}_{i}})$.
2.5 Fixed points and slant sums
Fix $\mathcal{M}^{(r)}:=\mathcal{M}_{Q^{(r)},\theta^{(r)}}(\mathsf{v}^{(r)},\mathsf{w}^{(r)})$ for $r=1,2$ as above, equipped with the actions for $\mathsf{T}^{(r)}$ for $r\in\{1,2\}$. Let $\star_{1}$ and $\star_{2}$ be compatible vertices. Let $\mathcal{M}:=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$, which is equipped with the action of a torus $\mathsf{T}$ where $\operatorname{rank}\mathsf{T}=\operatorname{rank}\mathsf{T}^{(1)}+\operatorname{rank}\mathsf{T}^{(2)}-\mathsf{w}^{(2)}_{\star_{2}}-1$.
Theorem 2.11.
Let $p^{(1)}$ be a fixed point on $\mathcal{M}^{(1)}$ split over $\star_{1}$. Choose a $(p^{(1)},\star_{1})$-chamber $\mathfrak{C}$ of $\mathsf{T}^{(1)}$. Assume that $\theta^{(2)}>0$ or $\theta^{(2)}<0$. Then there is a closed embedding
$\Psi:=\Psi_{p^{(1)},\mathfrak{C}}:(\mathcal{M}^{(2)})^{\mathsf{T}^{(2)}}\to\mathcal{M}^{\mathsf{T}}$
Proof.
Since $p^{(1)}$ is $\mathsf{T}^{(1)}$-fixed, there is a homomorphism $\rho:\mathsf{T}^{(1)}\to G_{\mathsf{v}^{(1)}}$ as in Section 2.1 which provides a decomposition of $V_{\star_{1}}$ into weight spaces of $\mathsf{T}^{(1)}$. Since $p^{(1)}$ is split over $\star_{1}$, the nonzero weight spaces are one-dimensional. The chamber $\mathfrak{C}$ totally orders the weight spaces. Making a choice of weight basis for $V_{\star_{1}}$, we obtain an isomorphism $\phi:V_{\star_{1}}\xrightarrow{\sim}W_{\star_{2}}$ which depends both on $\mathfrak{C}$ and the choice of basis.
Let $p^{(2)}\in(\mathcal{M}^{(2)})^{\mathsf{T}^{(2)}}$. Let $\pi_{r}:\mu_{r}^{-1}(0)^{s}\to\mathcal{M}^{(r)}$ be the quotient map for $r\in\{1,2\}$. Let $\pi_{1}^{\mathrm{inv}}$ be the restriction of $\pi_{1}$ to the locus of quiver representations which are invariant with respect to the action of $\mathsf{T}^{(1)}$ induced by the grading on $V_{i}$ for $i\in Q^{(1)}_{0}$ and the action on edges of the quiver. We will define a morphism
$\tilde{\Psi}:(\pi_{1}^{\mathrm{inv}})^{-1}(p^{(1)})\times\pi_{2}^{-1}(p^{(2)})\to\mathcal{M}$
which will descend to the map in the statement of the theorem. Note that the desired domain of $\tilde{\Psi}$ lies in the top left of (7), and we will use the notation $\bf{Y}$} and $\bf{Z}$} defined there. Our choice of isomorphism $\phi:V_{\star_{1}}\xrightarrow{\sim}W_{\star_{2}}$ discussed above provides a section $\pi_{1}^{-1}(p^{(1)})\times\pi_{2}^{-1}(p^{(2)})\to{\bf{Y}}$. Since $p^{(2)}$ is $\mathsf{T}^{(2)}$ fixed, Lemma 2.5 implies that the hypotheses of Lemma 2.8 hold, implying that this section factors through $\bf{Z}$}. Composing with the top right horizontal arrow followed by the rightmost vertical arrow in (7) provides the desired $\tilde{\Psi}$.
The map $\tilde{\Psi}$ is easily seen to be invariant under the action of $G_{\mathsf{v}^{(2)}}$ and factors of the centralizer $Z_{G_{\mathsf{v}^{(1)}}}(\rho(\mathsf{T}^{(1)}))$ corresponding to vertices besides $\star_{1}$. Since $p^{(1)}$ is split over $\star_{1}$, the factor of this centralizer in $GL(V_{\star_{1}})$ consists of matrices that are diagonal with respect to a choice of weight basis of $V_{\star_{1}}$. Using the isomorphism $\phi:V_{\star_{1}}\to W_{\star_{2}}$ and the fact that $p^{(2)}$ is $\mathsf{T}^{(2)}$-fixed, we can compensate for the action of these diagonal matrices by an element of $G_{\mathsf{v}^{(2)}}$. Thus $\tilde{\Psi}$ is also $Z_{G_{\mathsf{v}^{(1)}}}(\rho(\mathsf{T}^{(1)}))$ invariant and thus descends to a map $p^{(1)}\times p^{(2)}\to\mathcal{M}$. Varying $p^{(2)}$ provides the map $\Psi$ in the statement of the theorem.
We now show that $\Psi$ is independent of the choice of $\phi$, i.e. the choice of weight basis for $V_{\star_{1}}$. Let $\phi_{1},\phi_{2}:V_{\star_{1}}\to W_{\star_{2}}$ be two isomorphisms as in the first paragraph of this proof, leading to $\tilde{\Psi}_{1}$ and $\tilde{\Psi}_{2}$. Then $\tilde{\Psi}_{1}(x,y)=\tilde{\Psi}_{2}(x,t\cdot y)$, where $t:=\phi_{1}\circ\phi_{2}^{-1}\in\text{Aut}(W_{\star_{2}})\cap\mathsf{T}^{(2)}$. Since $p^{(2)}$ is $\mathsf{T}^{(2)}$-fixed, for any $y\in\pi_{2}^{-1}(p^{(2)})$, $\exists g\in G_{\mathsf{v}^{(2)}}$ such that $g\cdot y=t\cdot y$. So $\tilde{\Psi}_{2}(x,t\cdot y)=\tilde{\Psi}(x,g\cdot y)=(1,g)\cdot\tilde{\Psi}_{2}(x,y)$. In other words, $\tilde{\Psi}_{1}$ and $\tilde{\Psi}_{2}$ differ by an element of $G$ and thus descend to the same maps.
Finally we must justify why $\Psi$ lands in the $\mathsf{T}$-fixed locus. To show this, we define a inclusion
$\iota:\mathsf{T}=\mathsf{A}\times\mathbb{C}^{\times}_{\hbar}\hookrightarrow\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}=\mathsf{A}^{(1)}\times\mathbb{C}^{\times}_{\hbar^{(1)}}\times\mathsf{A}^{(2)}\times\mathbb{C}^{\times}_{\hbar^{(2)}}$ (8)
We have
  • a diagonal inclusion $\mathbb{C}^{\times}_{\hbar}\hookrightarrow\mathbb{C}^{\times}_{\hbar^{(1)}}\times\mathbb{C}^{\times}_{\hbar^{(2)}}$,
  • an inclusion
    $\mathsf{A}\times\mathbb{C}^{\times}_{\hbar}=\mathsf{A}^{(1)}\times\mathbb{C}^{\times}_{\hbar}\times\left(\prod_{\begin{subarray}{c}i\in Q^{(2)}_{0}\\ i\neq\star_{2}\end{subarray}}(\mathbb{C}^{\times})^{\mathsf{w}^{(2)}_{i}}\right)\subset\mathsf{A}^{(1)}\times\mathsf{A}^{(2)}$
    whose component inside of $\text{Aut}(W_{j})$ for $j\neq\star_{2}$ is induced by the identity maps and whose component inside $\text{Aut}(W_{\star_{2}})$ is given by combining the $\mathsf{T}^{(1)}=\mathsf{A}^{(1)}\times\mathbb{C}^{\times}_{\hbar}$ action on $V_{\star_{1}}$ with the isomorphism $\phi$
Putting these two maps together gives $\iota$, along which $\tilde{\Psi}$ is equivariant. So $\Psi$ is $\mathsf{T}$-fixed. ∎
The maps $\Psi_{p^{(1),\mathfrak{C}}}$ depend crucially on $p^{(1)}$. The dependence on $\mathfrak{C}$, however, is very mild.
Proposition 2.12.
Let $\mathfrak{C}$ and $\mathfrak{C}^{\prime}$ be $(p^{(1)},\star_{1})$-chambers. Let $F_{i}$, $i\in S$ be the fixed components of $\mathcal{M}^{(2)}$. There exist a permutation $s:S\to S$ and isomorphisms $f_{i}:F_{i}\to F_{s(i)}$ such that
$\Psi_{p^{(1)},\mathfrak{C}}\big|_{F_{i}}=\Psi_{p^{(1)},\mathfrak{C}^{\prime}}\big|_{F_{s(i)}}\circ f_{i}$
Proof.
The maps $\Psi_{p^{(1)},\mathfrak{C}}$ and $\Psi_{p^{(1)},\mathfrak{C}^{\prime}}$ depend on $\mathfrak{C}$ and $\mathfrak{C}^{\prime}$ only through the two isomorphisms $V_{\star_{1}}\xrightarrow{\sim}W_{\star_{2}}$. Composing one with the inverse of the other induces the desired $s$ and $f_{i}$. ∎
In other words, the maps $\Psi_{p^{(1)},\mathfrak{C}}$ depends on $\mathfrak{C}$ only up to permutation of the fixed components.
3 Quasimaps and slant sums
3.1 Review of quasimaps
As discussed in §2, Nakajima quiver varieties are defined as GIT quotients. so we may utilize the theory of quasimaps to a GIT quotient developed in [10]. In this paper, the domain of quasimaps will always be a parameterized $\mathbb{P}^{1}$.
Let $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ be a quiver variety, equipped with the action of $\mathsf{T}=\mathsf{A}\times\mathbb{C}^{\times}_{\hbar}$.
By definition a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}$ is a map $\mathbb{P}^{1}\to[\mu^{-1}(0)/G_{\mathsf{v}}]$. It is stable if it generically lands in $\mathcal{M}$, which is contained in the stack quotient as an open subset. The points for which the map does not land in $\mathcal{M}$ are called singularities of the quasimap.
So a quasimap consists of the data $(f,(\mathscr{V}_{i})_{i\in Q_{0}})$ where $\mathscr{V}_{i}$ is a vector bundle of rank $\mathsf{v}_{i}$ on $\mathbb{P}^{1}$ and $f\in H^{0}(\mathbb{P}^{1},\mathscr{N}\oplus\hbar^{-1}\otimes\mathscr{N}^{*})$ where
$\mathscr{N}=\bigoplus_{e\in Q_{1}}\operatorname{Hom}(\mathscr{V}_{t(e)},\mathscr{V}_{h(e)})\oplus\bigoplus_{i\in Q_{0}}\operatorname{Hom}(\mathscr{W}_{i},\mathscr{V}_{i})$
Here $\mathscr{W}_{i}$ denotes the trivial bundle $W_{i}\times\mathbb{P}^{1}$ on $\mathbb{P}^{1}$, which has a natural $\mathsf{A}$-equivariant structure, and $\hbar^{-1}$ denotes the trivial line bundle acted on by $\mathsf{T}$ with weight $\hbar^{-1}$.
The degree of a quasimap is defined to be $(\deg\mathscr{V}_{i})_{i\in Q_{0}}\in\mathbb{Z}^{Q_{0}}$.
Let $\mathsf{QM}$ be the moduli space of stable quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$. By [10], it is a Deligne-Mumford stack of finite type with a perfect deformation-obstruction theory. Thus it is equipped with a canonical virtual structure sheaf, which we will denote by $\mathcal{O}_{\mathrm{vir}}$. The canonical polarization $T^{1/2}$ of $\mathcal{M}$ pulls back under the universal evaluation morphism $\mathsf{QM}\times\mathbb{P}^{1}\to[\mu^{-1}(0)/G_{\mathsf{v}}]$, to a class $\mathscr{T}^{1/2}$, satisfying
$\mathscr{T}^{1/2}|_{(f,(\mathscr{V}_{i}))}=\bigoplus_{e\in Q_{1}}\operatorname{Hom}(\mathscr{V}_{t(e)},\mathscr{V}_{h(e)})\oplus\bigoplus_{i\in Q_{0}}\operatorname{Hom}(\mathscr{W}_{i},\mathscr{V}_{i})\ominus\bigoplus_{i\in Q_{0}}\operatorname{Hom}(\mathscr{V}_{i},\mathscr{V}_{i})$
The virtual tangent space at a quasimap $(f,(\mathscr{V}_{i}))$ is
$\mathscr{T}_{\mathrm{vir}}|_{(f,(\mathscr{V}_{i}))}=H^{*}\left(\mathbb{P}^{1},\mathscr{T}^{1/2}\oplus\hbar^{-1}\mathscr{T}^{1/2}\right)$ (9)
For $x\in\mathbb{P}^{1}$, let $\mathsf{QM}_{\mathrm{ns}\,x}$ be the moduli space of stable quasimaps which are nonsingular at $x$. The substack of degree $d$ quasimaps and the virtual sheaf on it will be denoted by $\mathsf{QM}^{d}_{\mathrm{ns}\,x}$ and $\mathcal{O}_{\mathrm{vir}}^{d}$ respectively.
We sometimes refer to an element of $\mathsf{QM}^{d}_{\mathrm{ns}\,x}$ only by its section $f$, though it also carries the information of the bundles $\mathscr{V}_{i}$. As stated, $\mathsf{QM}^{d}_{\mathrm{ns}\,x}$ depends explicitly on the presentation of $\mathcal{M}$ as a GIT quotient. In 3d mirror symmetry, this is viewed as a feature.
There is a natural action of $\mathbb{C}^{\times}$ on $\mathbb{P}^{1}$ such that $(\mathbb{P}^{1})^{\mathbb{C}^{\times}}=\{0,\infty\}$. We denote by $q$ the weight of $T_{0}\mathbb{P}^{1}$ and denote this rank one torus by $\mathbb{C}^{\times}_{q}$. The action of $\mathbb{C}^{\times}_{q}$ on $\mathbb{P}^{1}$ induces an action on $\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}$, as does the action of $\mathsf{T}$ on $\mathcal{M}$. There is a natural $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ equivariant structure on $\mathcal{O}_{\mathrm{vir}}^{d}$ and we will work $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-equivariantly.
There are morphisms
$\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}$$\left[\mu^{-1}(0)/G_{\mathsf{v}}\right]$$\mathcal{M}$$\scriptstyle{\operatorname{ev}_{0}}$$\scriptstyle{\operatorname{ev}_{\infty}}$
given by evaluating quasimaps at the respective points.
If a point $p\in\mathcal{M}$ is chosen, we define $\mathsf{QM}^{d}_{p}$ by the pullback diagram
$\mathsf{QM}^{d}_{p}$$\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}$$\{p\}$$\mathcal{M}$}
of quasimaps such that $f(\infty)=p$. The space $\mathsf{QM}^{d}_{p}$ is sometimes referred to as the moduli space of quasimaps which are based at $p$. If $p\in\mathcal{M}^{\mathsf{T}}$, then $\mathsf{QM}^{d}_{p}$ is preserved by the $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ action.
3.2 Twisted quasimaps
Choose a co character $\sigma:\mathbb{C}^{\times}\to\mathsf{A}$. We will also need to consider $\sigma$-twisted quasimaps to $\mathcal{M}$. The definition is identical to that of §3.1, except instead of the trivial bundles $\mathscr{W}_{i}$, we use the nontrivial bundles
$\mathscr{W}^{\sigma}_{i}:=(\mathbb{C}^{2}\setminus\{0\})\times W)/\mathbb{C}^{\times}$
where $\mathbb{C}^{\times}$ acts by scaling on the first factor and on the second by $\sigma$. Equivalently, a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}$ is a section of the trivial $\mu^{-1}(0)/G_{\mathsf{v}}]$ bundle on $\mathbb{P}^{1}$; a $\sigma$-twisted quasimap is a section of a nontrivial $\mu^{-1}(0)/G_{\mathsf{v}}]$ bundle on $\mathbb{P}^{1}$ whose topology is determined by $\sigma$.
Denote the moduli space of $\sigma$-twisted quasimaps by $\mathsf{QM}^{\sigma}$. All the discussion of §3.1 applies to $\sigma$-twisted quasimaps. In particular, we have $\mathsf{QM}^{\sigma,d}$, $\mathsf{QM}^{\sigma,d}_{\mathrm{ns}\,\infty}$, $\mathsf{QM}^{\sigma,d}_{p}$. The virtual tangent space is given by (9).
3.3 Vertex functions
The restriction of the evaluation map $\operatorname{ev}_{\infty}:\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}\to\mathcal{M}$ to the $\mathbb{C}^{\times}_{q}$-fixed locus is proper, see [37, §7.3]. So we obtain a pushforward in localized equivariant $K$-theory,
$(\operatorname{ev}_{\infty})_{*}:K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathsf{QM}^{d}_{\mathrm{ns}\,\infty})\to K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathcal{M})_{loc}:=K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathcal{M})\otimes\operatorname{Frac}K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(pt)$
The symmetrized virtual structure sheaf is defined by
$\hat{\mathcal{O}}_{\operatorname{vir}}=\mathcal{O}_{\mathrm{vir}}\otimes\left(\mathscr{K}_{\mathrm{vir}}\frac{\det\mathscr{T}^{1/2}|_{\infty}}{\det\mathscr{T}^{1/2}|_{0}}\right)^{1/2}$ (10)
where $\mathscr{K}_{\mathrm{vir}}=\operatorname{ev}_{\infty}^{*}(\det T\mathcal{M})\otimes\left(\det\mathscr{T}_{\mathrm{vir}}\right)^{-1}$ is the (normalized) virtual canonical bundle. The existence of the square root is discussed in [37, §6.1.8].
Enumerative counts of quasimaps are encoded in the generating function called the descendant vertex function. For a descendant $\tau\in K_{\mathsf{T}}([\mu^{-1}(0)/G_{\mathsf{v}}])$ this is defined as
$V^{(\tau)}(z)=\sum_{d\in\operatorname{Eff}(\mathcal{M})}(\operatorname{ev}_{\infty})_{*}\left(\mathsf{QM}^{d}_{\mathrm{ns}\,\infty},\operatorname{ev}_{0}^{*}(\tau)\otimes\hat{\mathcal{O}}_{\operatorname{vir}}\right)z^{d}$
Here the sum runs over the cone of effective quasimap classes, which is defined as the set of $d\in\mathbb{Z}^{Q_{0}}$ for which there exists a stable quasimap of degree $d$ and $z^{d}$ stands for the multidegree $\prod_{i\in Q_{0}}z_{i}^{d_{i}}$. Vertex functions are elements of
$K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathcal{M})_{loc}[[z]]:=\left\{\sum_{d}a_{d}z^{d}\,\mid\,d\in\operatorname{Eff}(\mathcal{M}),a_{d}\in K_{\mathsf{T}\times\mathbb{C}^{\times}_{q}}(\mathcal{M})_{loc}\right\}$
The case of $\tau=1$ is sometimes referred to as the bare vertex function and denoted $V(z):=V^{(1)}(z)$. For $p\in\mathcal{M}$, we denote the restriction by $V^{(\tau)}_{p}(z):=V^{(\tau)}(z)$.
Using twisted quasimaps, we can also define $\sigma$-twisted vertex functions, which we will denote by $V^{(\tau),\sigma}(z)$.
3.4 Branching rule for vertex functions
Now we return to slant sums and study vertex functions in the setting of Theorem 2.11, which we briefly recall.
Let $\mathcal{M}=\mathcal{M}^{(1)}\#\mathcal{M}^{(2)}$ be a slant sum over compatible vertices $\star_{1}$ and $\star_{2}$. We write the three quivers as $Q$, $Q^{(1)}$, and $Q^{(2)}$. Let $p^{(r)}\in(\mathcal{M}^{(r)})^{\mathsf{T}^{(r)}}$ and assume that $p^{(1)}$ is split over $\star_{1}$. Assume $\theta^{(2)}=\pm(1,1,\ldots,1)$. Choose a $(p^{(1)},\star_{1})$-chamber $\mathfrak{C}$ and let $p=\Psi_{p^{(1)},\mathfrak{C}}(p^{(2)})$ as in Theorem 2.11.
As in (8), we have an inclusion of tori $\iota:\mathsf{T}\hookrightarrow\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}$ such that $\mathcal{V}_{i}|_{p}=\iota^{*}\mathcal{V}^{(r)}_{i}|_{p^{(r)}}$ as $\mathsf{T}$-representations for $i\in Q^{(r)}$, $r\in\{1,2\}$. Furthermore, the composition $\mathsf{T}\xrightarrow{\iota}\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}\xrightarrow{pr_{2}}\mathsf{T}^{(2)}$ is surjective.
There are now several quiver varieties present and we will consider based quasimaps to each of them. The notation $\mathsf{QM}_{p}$ and $\mathsf{QM}_{p^{(1)}}$ now refer to quasimaps to the different varieties $\mathcal{M}$ and $\mathcal{M}^{(1)}$, respectively, rather than quasimaps to a fixed variety based at different points.
Any ordered tuple of integers $d=(d_{1},d_{2},\ldots,d_{n})$ where $n=\dim W_{\star_{2}}$ gives a cocharacter $\sigma_{d}$ of the framing torus $\mathsf{A}^{(2)}$, whose component acting on $W_{\star_{2}}$ is $(t)\mapsto\text{diag}(t^{d_{1}},t^{d_{2}},\ldots,t^{d_{n}})$ and is the identity on $W^{(2)}_{i}$ for $i\neq\star_{2}$.
We can obtain such a tuple from a $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-fixed quasimap $(f,(\mathscr{V}_{i}))$ in $\mathsf{QM}_{p}$ in the following way. Being fixed by $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ implies that each bundle $\mathscr{V}_{i}$ is endowed with a grading by $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{p}$. The $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{p}$ are exactly the same as the $\mathsf{T}^{(1)}$-weights of $\mathcal{V}^{(1)}_{i}|_{p^{(1)}}$ for $i\in Q^{(1)}$. In particular, $\mathscr{V}_{\star_{1}}$ is graded by distinct and ordered (by $\mathfrak{C}$) weights $w_{1},w_{2},\ldots,w_{n}$ of $\mathsf{T}^{(1)}$. Then we have $\mathscr{V}_{\star_{1}}\cong\bigoplus_{i=1}^{n}w_{i}\mathcal{O}(d_{i})$, giving rise to a tuple $(d_{1},d_{2},\ldots,d_{n})$.
By continuity, the cocharacter $\sigma_{f}$ depends only on the fixed component $F$ containing $(f,(\mathscr{V}_{i}))$. So we will also denote it by $\sigma_{F}$.
We are now ready to relate based quasimaps to the three quiver varieties.
Theorem 3.1.
There is an isomorphism
$\left(\mathsf{QM}^{d^{(1)},d^{(2)}}_{p}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}\cong\bigsqcup_{F}F\times\left(\mathsf{QM}^{\sigma_{F},d^{(2)}}_{p^{(2)}}\right)^{\mathsf{T}^{(2)}\times\mathbb{C}^{\times}_{q}}$
where $F$ runs over all connected components of $\left(\mathsf{QM}^{d^{(1)}}_{p^{(1)}}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$. Furthermore, this isomorphism identifies the virtual tangent spaces as $\mathsf{T}$, where the virtual tangent spaces on the right hand side are pulled back via $\iota$.
Proof.
By definition, a quasimap to $\mathcal{M}$ is a pair $(f,(\mathscr{V})_{i\in Q_{0}})$, where $\mathscr{V}_{i}$ are vector bundles on $\mathbb{P}^{1}$ and
$f\in H^{0}\left(\mathbb{P}^{1},\mathscr{N}\oplus\mathscr{N}^{*}\right)$
satisfying aforementioned conditions, where
$\mathscr{N}=\bigoplus_{e\in Q_{1}}\operatorname{Hom}(\mathscr{V}_{t(e)},\mathscr{V}_{h(e)})\oplus\bigoplus_{i\in Q_{0}}\operatorname{Hom}(\mathscr{W}_{i},\mathscr{V}_{i})$
and $\mathscr{W}_{i}$ is the trivial bundle of rank $\mathsf{w}_{i}$. This data is considered up to isomorphism of quasimaps, which are defined to be isomorphisms of the vector bundles (which are identities on $\mathscr{W}_{i}$).
Assume that the quasimap $f$ is $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-fixed and evaluates to $p$ at $\infty$. Being $\mathsf{T}$-fixed implies that each bundle $\mathscr{V}_{i}$ is graded by the $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{p}$. In particular, $\mathscr{V}_{\star_{1}}=\bigoplus_{j}w_{j}\mathscr{V}_{\star_{1},j}$ where $w_{j}$ run over distinct $\mathsf{T}$ weights, which are canonically ordered due to the choice of $\mathfrak{C}$. Let $\mathscr{W}_{\star_{2}}=\bigoplus_{j}a_{j}\mathcal{O}(d_{j})$ where $d_{j}=\deg\mathscr{V}_{\star_{1},j}$. Choose an isomorphism $\phi:\mathscr{V}_{\star_{1}}\to\mathscr{W}_{\star_{2}}$ equivariant with respect to $\iota$. The isomorphism $\phi$ determines and is determined by isomorphisms $\mathscr{V}_{\star_{1},j}\to\mathcal{O}(d_{j})$.
Being $\mathbb{C}^{\times}_{q}$-fixed implies that the quasimap evaluates to $p$ everywhere except possibly at $0\in\mathbb{P}^{1}$.
Let
$\mathscr{N}^{(1)}=\bigoplus_{e\in Q_{1}^{(1)}}\operatorname{Hom}(\mathscr{V}_{t(e)},\mathscr{V}_{h(e)})\oplus\bigoplus_{i\in Q_{0}^{(1)}}\operatorname{Hom}(\mathscr{W}_{i},\mathscr{V}_{i})$
and
$\mathscr{N}^{(2)}=\bigoplus_{e\in Q_{1}^{(2)}}\operatorname{Hom}(\mathscr{V}_{t(e)},\mathscr{V}_{h(e)})\oplus\bigoplus_{i\in Q_{0}^{(2)}}\operatorname{Hom}(\mathscr{W}_{i},\mathscr{V}_{i})\oplus\operatorname{Hom}(\mathscr{W}_{\star_{2}},\mathscr{V}_{\star_{2}})$
Let $f^{(1)}$ be the global section of $\mathscr{N}^{(1)}\oplus(\mathscr{N}^{(1)})^{*}$ determined by taking the respective components of $f$. The data $(f^{(1)},(\mathscr{V}_{i})_{i\in Q^{(1)}_{0}})$ defines a quasimap to $\mathcal{M}^{(1)}$. The quasimap $f^{(1)}$ is $\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}$-fixed and maps $\infty$ to $p^{(1)}$. Furthermore, an isomorphism $(f,(\mathscr{V}_{i})_{i\in Q_{0}})\cong(f^{\prime},(\mathscr{V}_{i})_{i\in Q_{0}})$ induces an isomorphism of $f^{(1)}$ and $f^{\prime(1)}$. Thus we have a well defined map
$\left(\mathsf{QM}_{p}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}\to\left(\mathsf{QM}_{p^{(1)}}\right)^{\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}}$
It is clear from the construction that if $\deg f=(d^{(1)},d^{(2)})$, then $\deg f^{(1)}=d^{(1)}$.
Using $\phi$, we similarly define a global section $f^{(2)}$ of $\mathscr{N}^{(2)}\oplus(\mathscr{N}^{(2)})^{*}$ by taking the relevant components of $f$.
We claim that the isomorphism class of $f^{(2)}$ does not depend on $\phi$. Let $\phi$ and $\phi^{\prime}$ be two such isomorphisms, leading to $f^{(2)}$ and $f^{\prime(2)}$. Note that $\phi^{\prime}\circ\phi^{-1}\in\text{Aut}(\mathscr{W}_{\star_{2}})$ preserves the $\mathsf{T}^{(2)}$-grading and is thus the same as the action of an element $t\in\mathsf{T}^{(2)}$. By definition, the quasimaps $f^{(2)}$ and $f^{\prime(2)}$ are related by $f^{(2)}=t\cdot f^{\prime(2)}$. Choose a lift $\tilde{t}\in\mathsf{T}$ of $t$ under the composition $\mathsf{T}\xrightarrow{\iota}\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}\to\mathsf{T}^{(2)}$. Since $f$ is $\mathsf{T}$-fixed, there is an isomorphism $\tilde{t}\cdot(f,(\mathscr{V}_{i})_{i\in Q_{0}})\cong(f,(\mathscr{V}_{i})_{i\in Q_{0}})$ which induces an isomorphism $(f^{(2)},(\mathscr{V}_{i})_{i\in Q^{(2)}_{0}})\cong(f^{\prime(2)},(\mathscr{V}_{i})_{i\in Q^{(2)}_{0}})$.
A similar argument shows that $(f,(\mathscr{V}_{i})_{i\in Q_{0}})\cong(f^{\prime},(\mathscr{V}^{\prime}_{i})_{i\in Q_{0}})$ implies that $(f^{(2)},(\mathscr{V}_{i})_{i\in Q_{0}^{(2)}})\cong(f^{\prime(2)},(\mathscr{V}^{\prime}_{i})_{i\in Q_{0}^{(2)}})$.
Overall, we obtain a degree preserving map
$\left(\mathsf{QM}_{p}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}\to\bigsqcup_{F}F\times\left(\mathsf{QM}_{p^{(2)}}^{\sigma_{F}}\right)^{\mathsf{T}^{(2)}\times\mathbb{C}^{\times}_{q}}$
The inverse is easy to construct.
It also follows immediately from (9) that the isomorphism of the theorem identifies the virtual tangent spaces as $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ representations, where we view the virtual tangent spaces on the right hand side as a $\mathsf{T}$ representation via the map $\iota:\mathsf{T}\hookrightarrow\mathsf{T}^{(1)}\times\mathsf{T}^{(2)}$.
Since it identifies the virtual tangent spaces, the isomorphism above respects the virtual structure sheaves. Since we assume the canonical polarization is used in (10), it follows that the isomorphism respects the symmetrized virtual structure sheaves as well. Recall the map $\iota$ from (1) and let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. We write the Kähler parameters similarly as $z=(z_{1},z_{2})$.
Theorem 3.2.
We have
$V^{(\tau)}_{p}(z_{1},z_{2})=\sum_{F}\chi\left(F,\frac{(\operatorname{ev}_{0}^{*}(\tau_{1})\otimes\hat{\mathcal{O}}_{\operatorname{vir}})|_{F}}{\bigwedge(N_{\operatorname{vir}}^{\vee}|_{F})}\right)z_{1}^{\deg F}\iota^{*}V^{(\tau_{2}),\sigma_{F}}_{p^{(2)}}(z_{2})$
where the sum runs over $\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}$-fixed components $F$ of $\mathsf{QM}_{p^{(1)}}$.
Proof.
This follows from Theorem 3.1 and the virtual localization theorem. ∎
For any quiver variety $\mathcal{M}$ and for any cocharacter $\sigma:\mathbb{C}^{\times}\to\mathsf{T}$, there is a bijection between torus fixed quasimaps and torus fixed twisted quasimaps. This bijection shifts the degree:
$\left(\mathsf{QM}^{d,\sigma}_{p}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}\cong\left(\mathsf{QM}^{\tilde{d}}_{p}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$
where $\tilde{d}_{i}=d_{i}+\langle\mathscr{L}_{i}|_{p},\sigma\rangle$ and $\langle\cdot,\cdot\rangle$ is the natural pairing between characters and cocharacters.
And the virtual tangent spaces at a torus fixed quasimap $f$ are related by
$\mathscr{T}_{\mathrm{vir}}^{\sigma}|_{f}-\mathscr{T}_{\mathrm{vir}}|_{f}=\frac{T^{\sigma}_{p}\mathcal{M}-T_{p}\mathcal{M}}{1-q}$ (11)
where $T^{\sigma}_{p}\mathcal{M}$ is the tangent space of $\mathcal{M}$ at $p$, viewed as a $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ representation where the second factor acts by $\sigma$. As in [37, §8.2], this leads to a relationship between the twisted and untwisted vertex functions.
To state it, it is convenient to define a few transcendental functions. Let $\Phi$ be the function defined on torus weights by $\Phi(x):=\prod_{i=0}^{\infty}(1-xq^{i})$ and extended to sums and differences of weights by multiplicativity. For a torus fixed point $p$ in a quiver variety, let $\Phi_{p}=\Phi((q-\hbar)T^{1/2}|_{p})$ and $\mathbf{e}_{p}=\exp\left(\ln(q)^{-1}\sum_{i}\ln(\mathscr{L}_{i}|_{p})\ln(z_{i})\right)$.
Proposition 3.3.
Let $\mathcal{M}$ be a quiver variety and let $p\in\mathcal{M}^{\mathsf{T}}$. Let $\sigma:\mathbb{C}^{\times}\to\mathsf{A}$ be a cocharacter. The twisted and untwisted vertex functions are related by
$V^{(\tau),\sigma}_{p}=\Phi_{p}^{-1}\mathbf{e}_{p}^{-1}\left(\Phi_{p}\mathbf{e}_{p}V^{(\tau)}_{p}\right)\big|_{a=aq^{\sigma}}$
Combining Proposition 3.3 and Theorem 3.2, we obtain the following.
Theorem 3.4.
With the notation as in Theorem 3.2, we have
$V^{(\tau)}_{p}(z_{1},z_{2})=\\ \iota^{*}\left(\mathbf{e}_{p^{(2)}}\Phi_{p^{(2)}}\right)^{-1}\sum_{F}\chi\left(F,\frac{(\operatorname{ev}_{0}^{*}(\tau_{1})\otimes\hat{\mathcal{O}}_{\operatorname{vir}})|_{F}}{\bigwedge(N_{\operatorname{vir}}^{\vee}|_{F})}\right)z_{1}^{\deg F}\iota^{*}\left(\left(\mathbf{e}_{p^{(2)}}\Phi_{p^{(2)}}V^{(\tau_{2})}_{p^{(2)}}\right)\big|_{a=aq^{\sigma_{F}}}\right)$
3.5 From branching to factorization
We explore here a few consequences of Theorem 3.4 when the vertex function of $\mathcal{M}^{(2)}$ is independent of the twist and thus can be factored out of the sum over $F$. We will freely use the notation of Theorem 3.4 in this section.
Corollary 3.5.
Assume that
  • $\tau_{2}|_{p^{(2)}}$ and $T\mathcal{M}^{(2)}|_{p^{(2)}}$ do not depend on the framing parameters of $W_{\star_{2}}$
  • $\mathcal{V}_{i}|_{p^{(2)}}$ is symmetric in the framing parameters of $W_{\star_{2}}$ for all $i\in Q^{(2)}_{0}$.
Then
$V^{(\tau_{1}\otimes\tau_{2})}_{p}(z_{1},z_{2})=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})V^{(\tau_{2})}_{p^{(2)}}(z_{2})$
where $z_{1,\star_{1}}^{\prime}=z_{1,\star_{1}}\prod_{i\in Q^{(2)}_{0}}z_{2,i}^{\deg_{\star_{2}}\mathscr{L}_{i}|_{p^{(2)}}}$ and $z_{1,j}^{\prime}=z_{1,j}$ otherwise.
See the proof for explanation of the notation $\deg_{\star_{2}}\mathscr{L}_{i}$.
Proof.
The first assumption combined with (11) implies that $V^{(\tau_{2}),\sigma}_{p^{(2)}}$ is independent of the twist $\sigma$, up to a monomial in the Kähler parameters. This monomial is accounted for by $\mathbf{e}_{p^{(2)}}$.
The second assumption implies that the degree of $\mathscr{L}_{i}|_{p^{(2)}}$ as a Laurent monomial in $a_{\star_{2},j}$ is independent of $j$. So we can denote it by $\deg_{\star_{2}}\mathscr{L}_{i}|_{p^{(2)}}$.
Let $F$ be a fixed component and let $d_{\star_{1}}\in\mathbb{Z}$ be the degree of the $\mathscr{V}_{\star_{1}}$ bundle on this component. Then
$\mathbf{e}_{p^{(2)}}|_{a=aq^{\sigma_{F}}}=\mathbf{e}_{p^{(2)}}\prod_{i\in Q^{(2)}_{0}}z_{2,i}^{d_{\star_{1}}\deg_{\star_{2}}\mathscr{L}_{i}|_{p^{(2)}}}$
Clearly, this monomial can be absorbed into a shift of $z_{1,\star_{1}}$ in the vertex function of $\mathcal{M}^{(1)}$. ∎
Remark 3.6.
One can show that the hypotheses of Corollary 3.5 are satisfied if $\mathcal{M}^{(2)}$ is zero dimensional and $\tau_{2}=1$, see also Proposition 4.2 below.
Due to form of the obstruction theory, the $q=\hbar$ specialization of vertex functions with descendant $1$ depends only on $\hbar$. In particular, it does not depend on any framing parameters.
Corollary 3.7.
Assume that
  • $\mathcal{V}_{i}|_{p^{(2)}}$ is symmetric in the framing parameters of $W_{\star_{2}}$ for all $i\in Q^{(2)}_{0}$
  • $\tau_{2}|_{p^{(2)}}$ is independent of the framing parameters of $W_{\star_{2}}$
Then
$V^{(\tau_{1}\otimes\tau_{2})}_{p}(z_{1},z_{2})\big|_{q=\hbar}=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})\big|_{q=\hbar}V^{(\tau_{2})}_{p^{(2)}}(z_{2})\big|_{q=\hbar}$
where $z_{1}^{\prime}$ is as in Corollary 3.5.
Another interesting special case occurs when $\mathsf{w}^{(2)}_{\star_{2}}=1$ and $\mathsf{w}^{(2)}_{i}=0$ otherwise, which most closely resembles the original slant sum constructions of [38].
Corollary 3.8.
Assume that $\mathsf{w}^{(2)}_{\star_{2}}=1$ and $\mathsf{w}^{(2)}_{i}=0$ otherwise. Then
$V^{(\tau_{1}\otimes\tau_{2})}_{p}(z_{1},z_{2})=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})V^{(\tau_{2})}_{p^{(2)}}(z_{2})$
where $z_{1,\star_{1}}^{\prime}=z_{1,\star_{1}}q^{\deg_{\star_{2}}\tau_{2}|_{p^{(2)}}}\prod_{i\in Q^{(2)}_{0}}z_{2,i}^{\deg_{\star_{2}}\mathscr{L}_{i}|_{p^{(2)}}}$ and $z_{1,j}^{\prime}=z_{1,j}$ otherwise.
Proof.
The proof is similar to Corollary 3.5, the only difference being the possible dependence of $\tau_{2}|_{p^{(2)}}$ on the framing parameter of $W_{\star_{2}}$. This is accounted for in the shift for $z_{1}^{\prime}$. ∎
4 Consequences
4.1 Factorization in zero-dimensional case
In this section, we will follow standard notation and denote simple roots by $\alpha_{i}$ and fundamental weights by $\varpi_{i}$. We will also denote $\kappa=q/\hbar$.
Corollary 3.5 provides a strategy to approach Conjecture 1.3. This conjecture is known only for $A$ and $D$ type quiver varieties with framings only at minuscule vertices, for which it was proven in [14] and [17] respectively.
For each of these cases, the difficult part is proving that the vertex function factorizes to a certain explicit product of $q$-binomial functions. Then it is a relatively straightforward exercise in root combinatorics to identify the terms in the factorization with certain roots as in Conjecture 1.3. This is demonstrated by the following example.
Example 4.1.
Consider the $A_{3}$ quiver variety $\mathcal{M}$ with $\mathsf{v}=(1,2,1)$, $\mathsf{w}=(0,1,0)$, and $\theta=(1,1,1)$. The vertex function (normalized as in (4)) can be computed by localization, which leads to the following formula:
$\mathsf{V}_{\mathcal{M}}$ $=\sum_{d_{i,j}}\left(\frac{q}{\hbar}\right)^{N(d)}z_{1}^{d_{1,1}}z_{2}^{d_{2,1}+d_{2,2}}z_{3}^{d_{3,1}}\frac{\left(\hbar\right)_{d_{2,1}}\left(\hbar^{2}\right)_{d_{2,2}}}{\left(q\right)_{d_{2,1}}\left(q\hbar\right)_{d_{2,2}}}\cdot$
$\frac{\left(\frac{q}{\hbar}\right)_{d_{2,1}-d_{2,2}}\left(q\hbar\right)_{d_{2,2}-d_{2,1}}}{\left(1\right)_{d_{2,1}-d_{2,2}}\left(\hbar^{2}\right)_{d_{2,2}-d_{2,1}}}\cdot\frac{\left(\hbar\right)_{d_{2,2}-d_{1,1}}\left(1\right)_{d_{2,1}-d_{1,1}}\left(\hbar\right)_{d_{3,1}-d_{2,1}}\left(1\right)_{d_{3,1}-d_{2,2}}}{\left(q\right)_{d_{2,2}-d_{1,1}}\left(\frac{q}{\hbar}\right)_{d_{2,1}-d_{1,1}}\left(q\right)_{d_{3,1}-d_{2,1}}\left(\frac{q}{\hbar}\right)_{d_{3,1}-d_{2,2}}}$
where the sum runs over the indices $d_{1,1}$, $d_{2,1}$, $d_{2,2}$, and $d_{3,1}$ that form a reverse plane partition over the Young diagram of $\lambda=(2,2)$ and $N(d)=-2d_{1,1}+d_{2,1}+d_{2,2}+2d_{3,1}$. Explicitly, the constraints are $0\leq d_{2,1}\leq d_{1,1}\leq d_{2,2}$ and $d_{2,1}\leq d_{3,1}\leq d_{2,2}$.
Although it is not obvious from the formula, the vertex function factorizes as
$\mathsf{V}_{\mathcal{M}}=\frac{\Phi\left(\hbar\kappa z_{2}\right)\Phi(\hbar z_{1}z_{2})\Phi(\hbar\kappa^{2}z_{2}z_{3})\Phi(\hbar\kappa z_{1}z_{2}z_{3})}{\Phi(\kappa z_{2})\Phi(z_{1}z_{2})\Phi(\kappa^{2}z_{2}z_{3})\Phi(\kappa z_{1}z_{2}z_{3})}$
The weights $\lambda$ and $\mu$ in this case are $\lambda=\varpi_{2}$ and $\mu=-\varpi_{2}$. It is easy to see that
$(\alpha,\mu)=\begin{cases}-1&\text{if $\alpha\in\{\alpha_{2},\alpha_{1}+\alpha_{2},\alpha_{2}+\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}\}$}\\ \geq 0&\text{otherwise}\end{cases}$
for any positive root $\alpha$. Recalling that $e^{\alpha_{1}}=\kappa^{-1}z_{1}$, $e^{\alpha_{2}}=\kappa z_{2}$, and $e^{\alpha_{3}}=\kappa z_{3}$, this confirms Conjecture 1.3 in this case.
Suppose that $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$ are zero-dimensional quiver varieties whose vertex functions are known to factorize into a product of $q$-binomial functions. By the discussion above, they could be type $A$ or $D$ quiver varieties with framings only at minuscule vertices. Suppose that $\star_{1}$ and $\star_{2}$ are compatible vertices, that $\mathcal{M}^{(1)}$ is split over $\star_{1}$, and that $\theta^{(2)}=\pm(1,1,\ldots,1)$. Consider $\mathcal{M}:=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$.
With these assumptions, Corollary 3.5 implies the following.
Proposition 4.2.
The vertex function of $\mathcal{M}$ factorizes into a product of $q$-binomial functions.
As remarked above, proving this factorization is difficult. Matching the terms with the roots in Conjecture 1.3 for $\mathcal{M}$ is now reduced to a combinatorial exercise.
Consider the following example.
Example 4.3.
Let $\mathcal{M}^{(1)}$ be the type $A_{3}$ quiver variety with $\mathsf{v}^{(1)}=(1,2,1)$ and $\mathsf{w}^{(1)}=(0,1,0)$. Let $\mathcal{M}^{(2)}$ be the type $A_{1}$ quiver variety with $\mathsf{v}^{(2)}=(2)$ and $\mathsf{w}^{(2)}=(2)$. Let $\star_{1}$ be the second vertex and let $\star_{2}$ be the first (and only) vertex. As a $\mathsf{T}^{(1)}$-representation, $\mathcal{V}_{\star_{1}}=1+\hbar$. So $\mathcal{M}^{(1)}$ is split over $\star_{1}$.
Let $\mathcal{M}=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$. It is a $D_{4}$ quiver variety with $\mathsf{v}=(1,2,1,2)$ and $\mathsf{w}=(0,1,0,0)$. Note that the framing is not minuscule, so this $D_{4}$ quiver variety is not treated in [14]. In agreement with our choice of labeling of vertices, we will use $z_{1},z_{2},z_{3}$ (resp. $z_{4}$) for the Kähler variables of $\mathcal{M}^{(1)}$ (resp. $\mathcal{M}^{(2)}$).
In pictures, we have
11221111#\#2222==1122112211
The vertex function of $\mathcal{M}^{(1)}$ was written in Example 4.1. Using localization, one sees that the vertex function of $\mathcal{M}^{(2)}$ is
$\mathsf{V}_{\mathcal{M}^{(2)}}=\sum_{d_{1},d_{2}\geq 0}\left(\kappa^{2}z_{4}\right)^{d_{1}+d_{2}}\left(\prod_{i,j=2}^{2}\frac{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}}}{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}}}\right)\left(\prod_{i,j=1}^{2}\frac{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}\right)$
see, for example [13]. It is not at all clear from the formula that $V_{\mathcal{M}^{(2)}}$ in independent of $a_{1}$ and $a_{2}$. Nevertheless, it was shown in [13] that
$\mathsf{V}_{\mathcal{M}^{(2)}}=\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
in agreement with Conjecture 1.3. This is equivalent to the statement that the Macdonald-Ruijsenaars operators of row type act diagonally on the Macdonald polynomial for the empty partition with a certain eigenvalue, see [34].
By Corollary 3.5 and accounting for the various normalizations, we obtain
$\mathsf{V}_{\mathcal{M}}=\frac{\Phi\left(\hbar\kappa^{2}z_{2}z_{4}\right)\Phi(\hbar\kappa z_{1}z_{2}z_{4})\Phi(\hbar\kappa^{3}z_{2}z_{3}z_{4})\Phi(\hbar\kappa^{2}z_{1}z_{2}z_{3}z_{4})}{\Phi(\kappa^{2}z_{2}z_{4})\Phi(z_{1}z_{2}z_{4}q/\hbar)\Phi(\kappa^{3}z_{2}z_{3}z_{4})\Phi(\kappa^{2}z_{1}z_{2}z_{3}z_{4})}\cdot\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
Associated to $\mathcal{M}$ are the weights $\lambda=\varpi_{2}$ and $\mu=\varpi_{2}-2\varpi_{4}=-\epsilon_{3}-\epsilon_{4}$. Using the explicit construction of the $D_{4}$ root system, it is easy to calculate that
$(\alpha,\mu)=\begin{cases}-2&\alpha=\alpha_{4}\\ -1&\alpha\in\{\alpha_{2}+\alpha_{4},\alpha_{1}+\alpha_{2}+\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}\\ \geq 0&\text{otherwise}\end{cases}$
for any positive root $\alpha$. So we again confirm Conjecture 1.3.
We provide one more example of our slant sum constructions where the corresponding Kac-Moody algebra is of indefinite type.
Example 4.4.
Let $\mathcal{M}^{(1)}$ be the type $D$ quiver variety with $\mathsf{v}^{(1)}=(1,1,1,2,3,2,2)$ and $\mathsf{w}^{(1)}=(0,0,\dots,1)$:
1111112233222211
This is an example of a minuscule framing. Let $\star_{1,1}$ be the unframed spin node and $\star_{1,2}$ be the framed spin node.
Let $\mathcal{M}^{(2)}$ have a single vertex and no loops, with $\mathsf{v}^{(2)}=(2),\mathsf{w}^{(2)}=(2)$, and let $\star_{2}$ be this single vertex. Let $\mathcal{M}^{(3)}$ be another copy of this same quiver variety and let $\star_{3}$ be the vertex.
Consider the quiver variety
$\mathcal{M}:=\left(\mathcal{M}^{(1)}{}_{\star_{1,1}}\#_{\star_{2}}\mathcal{M}^{(2)}\right){}_{\star_{1,2}}\#_{\star_{3}}\mathcal{M}^{(3)}$
This looks like
11111122332222222211
The corresponding Kac-Moody algebra is of indefinite type and is not hyperbolic, see [21], exercise 4.2. We will calculate the vertex function of $\mathcal{M}$ using Corollary 3.5.
Conjecture 1.3 claims that
$\mathsf{V}_{\mathcal{M}^{(1)}}=\prod_{m\in S^{(1)}}\frac{\Phi(\hbar m)}{\Phi(m)}$
where
$S^{(1)}=\{z_{1},\kappa^{2}z_{6},z_{1}z_{2},\kappa z_{5}z_{6},z_{4}z_{5}z_{6},\kappa^{3}z_{5}z_{6}z_{7},\kappa^{2}z_{4}z_{5}z_{6}z_{7},\kappa z_{4}z_{5}^{2}z_{6}z_{7},\\ \kappa^{-1}z_{1}z_{2}z_{3}z_{4}z_{5}z_{6},\kappa z_{1}z_{2}z_{3}z_{4}z_{5}z_{6}z_{7},z_{1}z_{2}z_{3}z_{4}z_{5}^{2}z_{6}z_{7},\kappa^{-1}z_{1}z_{2}z_{3}z_{4}^{2}z_{5}^{2}z_{6}z_{7}\}$
and $\kappa=q/\hbar$. This formula was proven in [14]. Using $z_{8}$ (resp. $z_{9}$) as the Kähler parameter of $\mathcal{M}^{(2)}$ (resp. $\mathcal{M}^{(3)}$), we have
$\mathsf{V}^{(2)}=\frac{\Phi(\hbar\kappa z_{8})\Phi(\hbar\kappa^{2}z_{8})}{\Phi(\kappa z_{8})\Phi(\kappa^{2}z_{8})},\quad\mathsf{V}^{(3)}=\frac{\Phi(\hbar\kappa z_{9})\Phi(\hbar\kappa^{2}z_{9})}{\Phi(\kappa z_{9})\Phi(\kappa^{2}z_{9})}$
By Corollary 3.5, we have
$\mathsf{V}_{\mathcal{M}}=\prod_{m\in S}\frac{\Phi(\hbar m)}{\Phi(m)}$
where
$S=\{z_{1},\kappa^{2}z_{6}z_{8},z_{1}z_{2},\kappa z_{5}z_{6}z_{8},z_{4}z_{5}z_{6}z_{8},\kappa^{3}z_{5}z_{6}z_{7}z_{8}z_{9},\kappa^{2}z_{4}z_{5}z_{6}z_{7}z_{8}z_{9},\kappa z_{4}z_{5}^{2}z_{6}z_{7}z_{8}z_{9},\\ \kappa^{-1}z_{1}z_{2}z_{3}z_{4}z_{5}z_{6}z_{8},\kappa z_{1}z_{2}z_{3}z_{4}z_{5}z_{6}z_{7}z_{8}z_{9},z_{1}z_{2}z_{3}z_{4}z_{5}^{2}z_{6}z_{7}z_{8}z_{9},\kappa^{-1}z_{1}z_{2}z_{3}z_{4}^{2}z_{5}^{2}z_{6}z_{7}z_{8}z_{9},\\ \kappa z_{8},\kappa^{2}z_{8},\kappa z_{9},\kappa^{2}z_{9}\}$
In the Kac-Moody algebra, the associated weights are $\lambda=\varpi_{7}$ and
$\mu=-\varpi_{1}+\varpi_{3}+\varpi_{6}+2\varpi_{7}-2\varpi_{8}-2\varpi_{9}$
We can observe directly that all roots $\alpha$ associated to the monomials of $S$ satisfy $(\alpha,\mu)<0$ and with the appropriate magnitude. Using Lemma 6.9, it is a straightforward exercise to check that these are the only positive, real roots satisfying this property.
4.2 Branching of the nonstationary Ruijsenaars function
For $n\geq 2$, let $X_{n}$ be the quiver variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ with $\mathsf{v}=(1,2,\dots,n-1)$ and $\mathsf{w}=(0,\dots,0,n)$, and $\theta=(1,1,\ldots,1)$ for the $A_{n-1}$ quiver:
[…]
It is known that $X_{n}\cong T^{*}\mathcal{F}\ell(\mathbb{C}^{n})$, the cotangent bundle to the variety parameterizing quotients $\mathbb{C}^{n}\twoheadrightarrow V_{n-1}\twoheadrightarrow\ldots\twoheadrightarrow V_{1}$ where $\dim V_{i}=i$. The slant sum construction can be used to relate $X_{n}$ and $X_{n-1}$.
Let $Y_{n-1}$ be the quiver variety $\mathcal{M}_{Q^{\prime},\theta^{\prime}}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})$ with $\mathsf{v}^{\prime}=(n-1)$ and $\mathsf{w}^{\prime}=(n)$ for the quiver with a single vertex and no edges:
n1n-1nn
with $\theta^{\prime}=(1)$. It is known that $Y_{n-1}$ is the cotangent bundle to the Grassmannian parameterizing quotients $\mathbb{C}^{n}\twoheadrightarrow V$ where $\dim V=n-1$.
Let $\star_{1}$ be the single vertex in the quiver corresponding to $Y_{n-1}$, and $\star_{2}$ be the rightmost vertex of the $A_{n-1}$ quiver. By definition, the slant sum over the compatible vertices $\star_{1}$ and $\star_{2}$ is
$X_{n}=Y_{n-1}\#X_{n-1}$
Informally, this equation says that the flag variety in $\mathbb{C}^{n}$ can be built by attaching a flag variety in $\mathbb{C}^{n-1}$ to a Grassmannian. This is not literally true, but the point of this paper is that this can be made sense of at the level of torus fixed points.
The torus $\mathsf{T}^{(1)}=(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{\times}_{\hbar}$ acts on $Y_{n-1}$. Let $a_{1},\ldots,a_{n}$ denote the equivariant parameters of the first factor. Let $\mathcal{W}_{n-1}$ and $\mathcal{V}_{n-1}$ denote the tautological bundles on $Y_{n-1}$ corresponding to the framing and gauge vertices, respectively. Let $p^{(1)}\in Y_{n-1}^{\mathsf{T}^{(1)}}$ be the torus fixed point such that
$\mathcal{V}_{n}\big|_{p^{(1)}}=a_{1}+\dots+a_{n-1}$
The torus $\mathsf{T}^{(2)}=(\mathbb{C}^{\times})^{n-1}\times\mathbb{C}^{\times}_{\hbar}$ acts on $X_{n-1}$. Let $b_{1},\ldots,b_{n-1}$ denote the equivariant parameters for the first factor. Let $\mathcal{W}_{n-2}$ and $\mathcal{V}_{i}$ for $1\leq i\leq n-2$ denote the tautological bundles on $X_{n-1}$. Let $p^{(2)}\in X_{n-1}^{\mathsf{T}^{(2)}}$ be the fixed point such that
$\mathcal{V}_{i}\big|_{p^{(2)}}=b_{1}+\dots+b_{i}$
We choose the chamber for $\mathsf{T}^{(1)}$ which orders the weights of $\mathcal{V}_{n-1}|_{p^{(1)}}$ as
$a_{1}<\dots<a_{n-1}$
Let $p=p^{(1)}\#p^{(2)}$.
Applying Theorem 3.4 with trivial descendant, we obtain a formula for the vertex function of $X_{n}$ at $p$.
Recall from [5] or [13] that there is a bijection
$\left(\mathsf{QM}_{p^{(1)}}^{d}\right)^{\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}}=\left\{(d_{1},\ldots,d_{n-1})\,\mid\,\sum_{i}d_{i}=d\right\}$
and for a tuple $f=(d_{1},\ldots,d_{n-1})$,
$\frac{\hat{\mathcal{O}}_{\operatorname{vir}}|_{f}}{\bigwedge(N_{\operatorname{vir}}|_{f}^{\vee})}z_{n-1}^{\deg f}=\left(-\frac{q}{\hbar^{1/2}}\right)^{n|f|}\left(\prod_{i=1}^{n-1}\prod_{j=1}^{n}\frac{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}}}{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}}}\right)\left(\prod_{i,j=1}^{n-1}\frac{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}\right)z_{n-1}^{|f|}$
We also calculate
$\mathbf{e}_{p^{(2)}}=\exp\left(\ln(q)^{-1}\sum_{i=1}^{n-2}\ln(b_{1}\ldots b_{i})\ln(z_{i})\right)$
and
$\Phi_{p^{(2)}}=\prod_{1\leq i<j\leq n-1}\frac{\Phi\left(q\frac{b_{j}}{b_{i}}\right)}{\Phi\left(\hbar\frac{b_{j}}{b_{i}}\right)}$
So
$\mathbf{e}_{p^{(2)}}|_{b_{i}=b_{i}q^{d_{i}}}=\mathbf{e}_{p^{(2)}}\prod_{i=1}^{n-2}z_{i}^{d_{1}+\ldots d_{i}}$
and
$\Phi_{p^{(2)}}|_{b_{i}=b_{i}q^{d_{i}}}=\Phi_{p^{(2)}}\prod_{1\leq i<j\leq n-1}\frac{\left(\hbar\frac{b_{j}}{b_{i}}\right)_{d_{j}-d_{i}}}{\left(q\frac{b_{j}}{b_{i}}\right)_{d_{j}-d_{i}}}$
Putting this all together, Theorem 3.4 gives
$V_{p}=\sum_{d_{1},\ldots,d_{n-1}\geq 0}\left(\prod_{i=1}^{n-1}\prod_{j=1}^{n}\frac{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}}}{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}}}\right)\left(\prod_{i,j=1}^{n-1}\frac{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}\right)\\ \left(\prod_{1\leq i<j\leq n-1}\frac{\left(\hbar\frac{b_{j}}{b_{i}}\right)_{d_{j}-d_{i}}}{\left(q\frac{b_{j}}{b_{i}}\right)_{d_{j}-d_{i}}}\right)\left(-\frac{q}{\hbar^{1/2}}\right)^{n|f|}\left(\prod_{i=1}^{n-1}z_{i}^{d_{1}+\ldots d_{i}}\right)V_{p^{(2)}}|_{b_{i}=a_{i}q^{d_{i}}}$
Up to changes in notation, this is exactly equation (3.11) from [35]. In fact, the similarity of (3.11) of [35] with the vertex function of the cotangent bundle of Grassmannian was part of this inspiration for this work.
5 Mirror symmetry of Higgs and Coulomb branches
In this section we recall the basic facts about Higgs and Coulomb branches that we will use.
Recall that $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ denotes a quiver variety corresponding to some quiver $Q=(Q_{0},Q_{1})$, $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$ and a generic stability parameter $\theta\in\mathbb{Z}^{Q_{0}}$. As we discussed in Section 2, there is a natural (projective) morphism:
$\pi\colon\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\to\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w}),\qquad$
where $\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w})=\mu^{-1}(0)/\!\!/G_{\mathsf{v}}$ is the Hamiltonian reduction of ${\bf{M}}=T^{*}{\bf N}$, where ${\bf N}=\operatorname{Rep}_{Q}(\mathsf{v},\mathsf{w})$, by the action of the group $G_{\mathsf{v}}$.
5.1 Higgs branches
The variety $\mathcal{M}_{Q,0}(\mathsf{v},\mathsf{w})$ is the Higgs branch of the quiver gauge theory corresponding to the pair $(G_{\mathsf{v}},{\bf{N}})$. More generally, an arbitrary pair $(G,{\bf{N}})$ of a reductive group $G$ and a finite dimensional representation $\bf{N}$} should define a certain gauge theory such that its Higgs branch $\mathcal{M}_{H}(G,{\bf{N}})$ is the Hamiltonian reduction $T^{*}{\bf{N}}/\!\!/\!\!/G$.
One important feature of quiver gauge theories (i.e. those $(G,{\bf{N}})$ that arise from a choice of a quiver $Q$ as above) is that we can consider the GIT quotient $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ depending on a stability parameter $\theta$. For generic $\theta$ (see Proposition 2.2) the variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is smooth, symplectic, and is a resolution of singularities of the image of $\pi\colon\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\rightarrow\mathcal{M}_{H}(G,{\bf{N}})$. So, whenever the map $\pi$ is surjective (see [11] for the criteria), we can extract the Higgs branch $\mathcal{M}_{H}(G,{\bf{N}})$ from the resolution $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$. However, in many cases (for example, if ${\mathsf{w}}=0$), $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is empty but $\mathcal{M}_{H}(G,{\bf{N}})$ is highly nontrivial. We will denote $\mathcal{M}_{H}(G,{\bf{N}})$ by $\mathcal{M}_{H}$ whenever $(G,{\bf{N}})$ are clear.
5.2 Coulomb branches
In [8], Braverman, Finkelberg, and Nakajima proposed a mathematical definition of the Coulomb branch $\mathcal{M}_{C}(G,{\bf{N}})$ associated to a gauge theory corresponding to a pair $(G,{\bf{N}})$ as above. The variety $\mathcal{M}_{C}(G,{\bf{N}})$ is an affine Poisson variety. Even for nice quiver gauge theories (e.g., when $Q$ is ADE and $\mathsf{v},\mathsf{w}$ are such that $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is nonempty), in contrast to $\mathcal{M}_{H}(G,{\bf{N}})$, the varieties $\mathcal{M}_{C}(G,{\bf{N}})$ do not admit symplectic resolutions.
On the other hand, it is known that $\mathcal{M}_{C}(G,{\bf{N}})$ always has symplectic singularities, see [2]. In particular, $\mathcal{M}_{C}(G,{\bf{N}})$ always admits a $\mathbb{Q}$-factorial terminalization $\widetilde{\mathcal{M}}_{C}(G,{\bf{N}})$, which is a partial resolution of $\mathcal{M}_{C}(G,{\bf{N}})$ with nice properties (see [28, Section 4] for the general discussion). We will denote $\mathcal{M}_{C}(G,{\bf{N}})$ by $\mathcal{M}_{C}$ whenever $(G,{\bf{N}})$ are clear.
Let us briefly recall the definition of $\mathcal{M}_{C}$. Set $\mathcal{K}:=\mathbb{C}((z))$, $\mathcal{O}:=\mathbb{C}[[z]]$ and let $G_{\mathcal{K}}$, $G_{\mathcal{O}}$, ${\bf{N}}_{\mathcal{O}}$ be the corresponding spaces whose $\mathbb{C}$-points are $G((z))$, $G[[z]]$, and ${\bf{N}}[[z]]$ respectively. Set
$\mathcal{T}:=G_{\mathcal{K}}\times^{G_{\mathcal{O}}}{\bf{N}}_{\mathcal{O}},~\mathcal{R}=\{[g,n]\in\mathcal{T}\,|\,gn\in\mathbf{N}_{\mathcal{O}}\}.$
The group $G_{\mathcal{O}}$ acts naturally on the space $\mathcal{R}$. In [8, 2(ii)], the authors define the equivariant Borel-Moore homology $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ of $\mathcal{R}$. Moreover, in [8, Section 3], they endow $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ with the algebra structure given by convolution *. Finally, they prove that * is commutative and define:
$\mathcal{M}_{C}:=\operatorname{Spec}(H_{*}^{G_{\mathcal{O}}}(\mathcal{R}),*).$
One can also define a quantized Coulomb branch $\mathcal{A}_{C,\hbar}$ by adding an additional $\mathbb{C}^{\times}$-equivariance with respect to the “loop rotation” action:
$\mathcal{A}_{C,\hbar}:=H_{*}^{G_{\mathcal{O}}\rtimes\mathbb{C}^{\times}}(\mathcal{R}),~\mathcal{A}_{C}:=\mathcal{A}_{C,\hbar}|_{\hbar=1}.$
Convolution equips $\mathcal{A}_{C,\hbar}$ with an associative algebra structure.
5.3 Torus action and fixed points on Coulomb branches
The space $\mathcal{R}$ maps naturally to the affine Grassmannian $\operatorname{Gr}_{G}=G_{\mathcal{K}}/G_{\mathcal{O}}$. Connected components of $\operatorname{Gr}_{G}$ are labeled by the lattice $\pi_{1}(G)$. So, $\mathbb{C}[\mathcal{M}_{C}]$ admits a grading by $\pi_{1}(G)$, and hence, the variety $\mathcal{M}_{C}$ admits a (Hamiltonian) action of the torus $\pi_{1}(G)^{\wedge}$, where by $\bullet^{\wedge}$ we mean the Pontryagin dual (see [8, Section 3(v)] for the details). The cohomological grading on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ defines the action of $\mathbb{C}^{\times}$ on $\mathcal{M}_{C}$.
A choice of the stability parameter $\theta\colon G\rightarrow\mathbb{C}^{\times}$ defines a cocharacter $\mathbb{C}^{\times}\rightarrow\pi_{1}(G)^{\wedge}$. Abusing notations, we will denote this cocharacter by $\theta$. In [6, Conjecture 3.25(1)] the authors conjecture that the set $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}(\mathbb{C})$ of $\mathbb{C}$-points of the schematic fixed points of $\mathcal{M}_{C}$ is either empty or consists of one point. From now on, we denote $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}$ simply by $\mathcal{M}_{C}^{\theta}$.
5.4 Mirror symmetry and some conjectures
In this section we formulate various conjectures relating Higgs and Coulomb branches for quiver theories.
3d mirror symmetry is a nontrivial relation between Higgs and Coulomb branches for $(G,{\bf{N}})$. From now on, we assume that our theory is a quiver gauge theory for a quiver $Q$ without loops, and we denote the smooth symplectic variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ by $\widetilde{\mathcal{M}}_{H}$ (recall that in many cases $\widetilde{\mathcal{M}}_{H}$ is a symplectic resolution of the Higgs branch $\mathcal{M}_{H}$).
One mirror symmetry statement is the Hikita conjecture [20] predicting an identification of (graded) algebras:
$H^{*}(\widetilde{\mathcal{M}}_{H})\simeq\mathbb{C}[\mathcal{M}_{C}^{\theta}].$ (12)
So, it is natural to expect that $\mathcal{M}_{C}^{\theta}$ is nonempty if and only if $\widetilde{\mathcal{M}}_{H}$ is nonempty.
We start with the following standard lemma.
Lemma 5.1.
Let $\pi\colon X\rightarrow Y$ be a conical symplectic resolution. Assume that $H^{*}(X)=\mathbb{C}$. Then $X$ is isomorphic to $\mathbb{A}^{2k}$, where $2k=\operatorname{dim}X$.
Proof.
Let $p\in Y$ be the unique $\mathbb{C}^{\times}$-fixed point of $Y$ (recall that $Y$ contracts to $p$ via the action of $\mathbb{C}^{\times}$). The $\mathbb{C}^{\times}$-action contracts $X$ to $X^{\mathbb{C}^{\times}}=\pi^{-1}(p)^{\mathbb{C}^{\times}}$, so $H^{*}(X)=H^{*}(\pi^{-1}(p)^{\mathbb{C}^{\times}})$. Note now that $\pi^{-1}(p)^{\mathbb{C}^{\times}}$ is smooth and proper, so $H^{\mathrm{top}}(\pi^{-1}(p)^{\mathbb{C}^{\times}})=\mathbb{C}$. It follows that $\pi^{-1}(p)$ consists of one point. We conclude that $X$ contracts to one point under the action of $\mathbb{C}^{\times}$, hence, by the Bialynicki-Birula theorem (see [4]), $X$ must be isomorphic to the affine space. ∎
We make the following conjecture:
Conjecture 5.2.
The variety $\widetilde{\mathcal{M}}_{H}$ is a point if and only if $\mathcal{M}_{C}^{\theta}$ consists of a single nonsingular point.
Let us explain our motivation for this conjecture. If $p^{!}\in\mathcal{M}_{C}$ is nonsingular, then $\mathbb{C}[\mathcal{M}_{C}^{\theta}]$ must be equal to $\mathbb{C}$. Assuming (12), we deduce that $H^{*}(\widetilde{\mathcal{M}}_{H})=\mathbb{C}$. Then Lemma 5.1 implies that $\widetilde{\mathcal{M}}_{H}$ is isomorphic to $\mathbb{A}^{2k}$. We expect that for quivers $Q$ without edge loops, the dimension of $\widetilde{\mathcal{M}}_{H}$ is actually equal to zero whenever $p^{!}\in\mathcal{M}_{C}$ is nonsingular.
Recall that Conjecture 1.5 gives a formula for the tangent space $T_{p^{!}}\mathcal{M}_{C}$ when $\widetilde{\mathcal{M}}_{H}$ is a point. Actually we expect more to be true: this formula should allow one to compute the character of the tangent space for an arbitrary nonsingular fixed point of a partially resolved Coulomb branch. Let us explain the details.
Let $\mathsf{A}$ be a torus acting on $\bf{N}$} commuting with $G$ and let $\nu\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}$ be a cocharacter. As in [8, Section 3(ix)], it defines a partial resolution
$\widetilde{\mathcal{M}}_{C,\nu}\rightarrow\mathcal{M}_{C}.$ (13)
The action of $\pi_{1}(G)^{\vee}$ extends to $\widetilde{\mathcal{M}}_{C,\nu}$ and the morphism (13) is $\pi_{1}(G)^{\vee}$-equivariant. On the Higgs side, cocharacter $\nu$ defines an action $\mathbb{C}^{\times}\curvearrowright\widetilde{\mathcal{M}}_{H}$.
One can show that:
$\widetilde{\mathcal{M}}_{H}(G,{\bf{N}})^{\nu}=\bigsqcup_{\gamma\colon\mathbb{C}^{\times}\rightarrow G}\widetilde{\mathcal{M}}_{H}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)}),$
where the disjoint union is taken over all conjugacy classes of cocharacters $\gamma\colon\mathbb{C}^{\times}\rightarrow G$. Moreover, every $(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ corresponds to some quiver, hence, each $\widetilde{\mathcal{M}}_{H}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ itself is a Nakajima quiver variety.
Assume now that $\widetilde{\mathcal{M}}_{H}^{\nu}$ contains an isolated fixed point $p$. By the above, this point itself is equal to the Nakajima quiver variety corresponding to some choice of $\gamma$. Assuming Conjecture 5.2 holds, we see that the Coulomb branch $\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ has a unique nonsingular $\theta$-fixed point $p^{!}$.
The following conjecture is due to Justin Hilburn (see [23, Conjecture 1.6]).
Conjecture 5.3.
There is a $\pi_{1}(G)^{\wedge}$-equivariant open embedding:
$j\colon\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})\hookrightarrow\widetilde{\mathcal{M}}_{C,\nu}.$
for each $\gamma$.
Assuming both Conjectures 5.2 and 5.3 hold, one obtains the following corollary.
Corollary 5.4.
The point $j(p^{!})\in\widetilde{\mathcal{M}}_{C,\nu}$ is nonsingular and $j$ induces an isomorphism of tangent spaces at $p^{!}$ and $j(p^{!})$.
Recall now that Conjecture 1.5 describes the tangent space $T_{p^{!}}\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$. So, combining it with Conjecture 5.3, we should obtain a conjectural description of the tangent space to an nonsingular $\theta$-fixed point of $\widetilde{\mathcal{M}}_{C,\nu}$.
Remark 5.5.
We expect that there should be a bijection between nonsingular $\theta$-fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$ and isolated $\nu$-fixed points on $\widetilde{\mathcal{M}}_{H}$. If that is the case, then a tangent space to arbitrary non-singular $\theta$-fixed point should have a description as above. Note that, more generally, it is reasonable to expect that there should be a bijection between the set $\widetilde{\mathcal{M}}_{C,\nu}^{\theta}(\mathbb{C})$ and the set $\operatorname{Comp}(\widetilde{\mathcal{M}}_{H}^{\nu})$ of components of the $\nu$-fixed points of $\widetilde{\mathcal{M}}_{H}^{\nu}$ (this is another incarnation of the equivariant Hikita conjecture).
Remark 5.6.
Note that we are not assuming that the torus $\mathsf{A}$ comes from the action of $\prod_{i}GL(W_{i})$ on $\bf{N}$}. If this is the case, then $\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ is the product of Coulomb branches for the same quiver $Q$.
Let us finish this section by mentioning that whenever Conjecture 1.5 is proven, we obtain an explicit formula for a normalized character of the (unique) irreducible module in the category $\mathcal{O}$ for $\mathcal{A}$. Namely, assume that Conjecture 1.5 holds for some $Q,\mathsf{v},\mathsf{w}$. Fix $x\in\mathfrak{a}$ and let
$\mathcal{A}_{x}:=H_{*}^{(G_{\mathcal{O}}\rtimes\mathbb{C}^{\times})\times\mathsf{A}}(\mathcal{R})_{(x,1)}$
be the corresponding quantized Coulomb branch algebra. The cocharacter $\theta$ determines the category $\mathcal{O}_{\theta}(\mathcal{A}_{x})$ as in [27, Section 3.4]. By, [27, Lemma 3.27], irreducible objects of this category are in bijection with irreducible modules over the (finite-dimensional) algebra $\mathsf{C}_{\theta}(\mathcal{A}_{x})$ called Cartan subquotient or $B$-algebra and defined as follows:
$\mathsf{C}_{\theta}(\mathcal{A}_{x}):=\mathcal{A}_{x,0}/\sum_{i<0}\mathcal{A}_{x,i}\mathcal{A}_{x,-i},$
where index $i$ corresponds to the $\mathbb{Z}$-grading on $\mathcal{A}_{x}$ induced by the cocharacter $\theta$. It is known (see [27, Lemma 3.24]) that there exists a filtration on $\mathsf{C}_{\theta}(\mathcal{A}_{x})$ such that the associated graded $\operatorname{gr}\mathsf{C}_{\theta}(\mathcal{A}_{x})$ is a quotient of $\mathbb{C}[\mathcal{M}_{C}^{\theta}]$. Note now that $\mathbb{C}[\mathcal{M}_{C}^{\theta}]\simeq\mathbb{C}$, since $p^{!}=\mathcal{M}_{C}^{\theta}(\mathbb{C})$ is a nonsingular point of $\mathcal{M}_{C}$. So, we conclude that the category $\mathcal{O}_{\theta}(\mathcal{A}_{x})$ has a unique irreducible object $L=L(p^{!})$. Let $\widetilde{\operatorname{ch}}(L)$ be the $\mathsf{A}_{Q}$-character of $L$ normalized in such a way that it starts with $1$.
Conjecture 5.7.
For $\theta>0$ we have:
$\widetilde{\operatorname{ch}}(L)=\prod_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\frac{1}{(1-e^{\alpha})^{\langle\alpha,\mu\rangle}}.$ (14)
Remark 5.8.
One should be able to prove this conjecture by using microlocalization of $\mathcal{A}_{x}$ to a sheaf of algebras on $\mathcal{M}_{C}$. Since $\mathcal{M}_{C}$ is not conical, one should be careful. We will return to this question in the second version of the paper.
Remark 5.9.
Finally, let’s note that the quantized counterpart of Conjecture 5.3 should claim that there exists a homomorphism between the corresponding quantized Coulomb branches. Then, taking modules $L$ as above, one would obtain a collection of modules over the quantization of $\mathcal{M}_{C}$ labeled by nonsingular fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$. Let’s decompose $\lambda=\lambda_{1}+\ldots+\lambda_{N}$ into the sum of fundamental weights. Then, we obtain the decomposition $\mu=\mu_{1}+\ldots+\mu_{N}$ ($\mu_{i}=w(\lambda_{i})$). The module $L$ should be isomorphic to a tensor product of modules $L_{i}$ corresponding to $\lambda_{i},\mu_{i}$. For $Q$ of type ADE the corresponding modules $L_{i}$ are so-called prefundamental modules over $Y_{\mu_{i}}$.
6 Coulomb branches for ADE quivers
The goal of this section is to prove Conjectures 1.5, 5.2, and 5.3 for ADE quiver theories. Conjectures 5.2 and 5.3 can be deduced from the existing literature and our argument for the proof of Conjecture 1.5 is based on the approach used in [25]. From now on, we assume that $Q$ is of type ADE.
6.1 Realization via slices in affine Grassmannians
In this section, we recall all of the objects we need to prove conjectures above. Let $\mathfrak{g}=\mathfrak{g}_{Q}$ be the simple Lie algebra corresponding to $Q$. Let $G=G_{Q}$ be the adjoint Lie group with Lie algebra $\mathfrak{g}$ and choose a Borel subalgebra $B$ with opposite Borel $B_{-}$. Let $\mathsf{A}_{Q}$ be the maximal torus of $G$ (i.e. the intersection of $B$ and $B_{-}$) and let $\mathfrak{a}_{Q}$ be its Lie algebra. For $i\in Q_{0}$ we will denote by $\omega_{i}^{\vee},\alpha_{i}^{\vee}$ the corresponding fundamental coweight and simple coroot for $\mathfrak{g}$. Given $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$, set:
$\lambda=\sum_{i}\mathsf{w}_{i}\omega_{i}^{\vee},~\mu=\lambda-\sum_{i}\mathsf{v}_{i}\alpha_{i}^{\vee}.$
Following [6], we define the generalized transversal slice in the affine Grassmannian $\overline{\mathcal{W}}_{\mu}^{\lambda}$. It is the moduli space of the data $(\mathcal{P},\sigma,\phi)$, where:
  • $\mathcal{P}$ is a $G$-bundle on $\mathbb{P}^{1};$
[…]
  • $\phi$ is a $B$-structure on $\mathcal{P}$ (i.e. a $B$-subbundle of $\mathcal{P}$) of degree $w_{0}(\mu)$, having fiber $B_{-}$ at $\infty$ (with respect to $\sigma$).
The open subvariety $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$ consists of triples $(\mathcal{P},\sigma,\phi)$ above such that the degree of a pole of $\sigma$ at zero is equal to $\lambda$.
It follows from [6, Section 2(xi)] that $\overline{\mathcal{W}}^{\lambda}_{\mu}$ has the following matrix description:
$\overline{\mathcal{W}}^{\lambda}_{\mu}\simeq U[[z^{-1}]]_{1}\mathsf{A}_{Q}[[z^{-1}]]_{1}z^{\mu}U_{-}[[z^{-1}]]_{1}\cap\overline{(G[z]z^{\lambda}G[z])}.$ (15)
where the notation is as in [6]. The open subvariety $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$ consists of points $g\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lie in $G[z]z^{\lambda}G[z]$.
It follows from [6, Theorem 3.10] that we have an isomorphism:
$\mathcal{M}_{C}(\mathsf{v},\mathsf{w})\simeq\overline{\mathcal{W}}_{\mu}^{\lambda}.$ (16)
Remark 6.1.
For a coweight $\eta\in\mathfrak{a}_{Q}$ of $\mathfrak{g}$ we will denote by $\eta^{*}$ the coweight $w_{0}(\eta)$ (here $w_{0}$ is the longest element of the Weyl group $W$ acting naturally on $\mathfrak{a}_{Q}$). Strictly speaking, in [6, Theorem 3.10] the author construct an isomorphism $\mathcal{M}_{C}(\mathsf{v},\mathsf{w})\simeq\overline{\mathcal{W}}_{\mu^{*}}^{\lambda^{*}}$. Note now that there exists an identification:
$\overline{\mathcal{W}}^{\lambda}_{\mu}\simeq\overline{\mathcal{W}}^{\lambda^{*}}_{\mu^{*}}.$ (17)
[…]
The torus $\pi_{1}(G)^{\wedge}$ identifies with the maximal torus $\mathsf{A}_{Q}\subset G_{Q}$ acting naturally on $\overline{\mathcal{W}}_{\mu}^{\lambda}$ (changing the trivialization $\sigma$ at $\infty$). The action of $\mathbb{C}^{\times}_{\hbar}$ corresponds to the loop rotation action on $\overline{\mathcal{W}}_{\mu}^{\lambda}$ via the automorphisms of the curve $\mathbb{P}^{1}$.
Let ${\mathsf{A}}\subset\prod_{i}GL_{\mathsf{w}_{i}}$ be the flavor torus (i.e. the framing torus on the quiver variety side). Pick a cocharacter $\nu\colon\mathbb{C}^{\times}\rightarrow{\mathsf{A}}$. Let $N\in\mathbb{Z}_{\geqslant 1}$ be the number of eigenvalues of $\nu$. The choice of $\nu$ corresponds to the decomposition:
$\mathsf{w}_{i}=\mathsf{w}_{i}^{(1)}+\ldots+\mathsf{w}_{i}^{(N)},~i\in Q_{0}.$
Recall that we can associate to $\nu$ the partial resolution $\widetilde{\mathcal{M}}_{C,\nu}$ of $\mathcal{M}_{C}$. Let us recall the description of $\widetilde{\mathcal{M}}_{C,\nu}$ in the language of affine Grassmannian slices.
Consider the $N$-tuple of dominant coweights $\underline{\lambda}:=(\lambda_{1},\ldots,\lambda_{N})$ of $G_{Q}$, where
$\lambda_{j}=\sum_{i\in Q_{0}}\mathsf{w}_{i}^{(j)}\omega_{i}^{\vee}.$
Clearly, $\lambda=\lambda_{1}+\ldots+\lambda_{N}$. Set $\mathsf{T}_{Q}:={\mathsf{A}}_{Q}\times\mathbb{C}^{\times}_{\hbar}$. It follows from [7, Section 5] that we have a $\mathsf{T}_{Q}$-equivariant isomorphism of varieties:
$\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}\simeq\widetilde{\mathcal{M}}_{C,\nu},$ (18)
where $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ is the convolution diagram over $\overline{\mathcal{W}}^{\lambda}_{\mu}$ defined as follows.
The variety $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ is the moduli space of the data
$(\mathcal{P}^{\mathrm{triv}}=\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{N},\sigma_{1},\ldots,\sigma_{N},\phi),$
where
  • $\mathcal{P}_{i}$ is a $G$-bundle on $\mathbb{P}^{1}$;
[…]
  • $\phi$ is a $B$-structure on $\mathcal{P}_{N}$ of degree $w_{0}\mu$, having fiber $B_{-}$ at $\infty$ with respect to $\sigma_{N}\circ\sigma_{N-1}\circ\ldots\circ\sigma_{1}$.
We have a natural (proper and birational) morphism $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}\rightarrow\overline{\mathcal{W}}^{\lambda}_{\mu}$ given by
$(\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{N},\sigma_{1},\ldots,\sigma_{N},\phi)\mapsto(\mathcal{P}_{N},\sigma_{N}\circ\sigma_{N-1}\circ\ldots\circ\sigma_{1},\phi).$
Choose generic $\theta$. We have the following lemma.
Lemma 6.2.
The set of $\theta$-fixed points $(\overline{\mathcal{W}}^{\lambda}_{\mu})^{\theta}$ consists of one point if $\mu$ is a weight of $V(\lambda)$, the irreducible representation of the Langlands dual of $\mathfrak{g}$ of highest weight $\lambda$, and is empty otherwise. We will denote this fixed point by $z^{\mu}$.
Proof.
See [24, Lemma 2.8]. ∎
More generally, it follows from [25, Proposition 5.9] that the $\theta$-fixed points on $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ are isolated and are in bijection with $N$-tuples of coweights $\underline{\mu}=(\mu_{1},\ldots,\mu_{N})$ such that $\mu=\mu_{1}+\ldots+\mu_{N}$ and $\mu_{i}$ is a weight of $V(\lambda_{i})$. We will denote the point corresponding to $\underline{\mu}$ by $z^{\underline{\mu}}$. In [6, 2(vi)], the authors defined the so-called multiplication morphism: ${\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}\colon\overline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times\ldots\times\overline{\mathcal{W}}^{\lambda_{N}}_{\mu_{N}}\rightarrow\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lifts to a morphism:
$\tilde{{\bf{m}}}^{\underline{\lambda}}_{\underline{\mu}}\colon\overline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times\ldots\times\overline{\mathcal{W}}^{\lambda_{N}}_{\mu_{N}}\rightarrow\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}.$ (19)
The morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is known to be an open embedding (see [25, Proposition 5.7]). It is also known to be ${\mathsf{A}}_{Q}$-equivariant, and moreover, it becomes $\mathbb{C}^{\times}_{\hbar}$-equivariant after appropriately twisting the $\mathbb{C}^{\times}_{\hbar}$-action on the source of (19) (see [25, Section 5.4] for the details). It sends the unique $\mathsf{A}_{Q}$-fixed point of the domain to the point $z^{\underline{\mu}}\in\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$.
We also have a natural morphism $p\colon\overline{\mathcal{W}}^{\lambda}_{\mu}\rightarrow\overline{\operatorname{Gr}}^{\lambda}$ given by forgetting the $B$-structure $\phi$.
6.2 Proofs for ADE quivers
From now on, we assume that $\overline{\mathcal{W}}^{\lambda}_{\mu}$ contains a $\theta$-fixed point, i.e., that $\mu$ is a weight of $V(\lambda)$. Note that this precisely corresponds to $\widetilde{\mathcal{M}}_{H}:=\widetilde{\mathcal{M}}_{H}({\mathsf{v}},{\mathsf{w}})$ being nonempty (compare with the proof of Lemma 6.4 below).
Lemma 6.3.
The unique $\theta$-fixed point $z^{\mu}\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ is nonsingular iff $\mu\in W\lambda$.
Proof.
It follows from [29, Theorem 1.2] combined with [6, Remark 3.19] that the regular locus of $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is equal to $\mathcal{W}^{\lambda}_{\mu}$. It remains to note that $z^{\mu}\in\mathcal{W}^{\lambda}_{\mu}$ iff $\mu\in W\lambda$. ∎
Lemma 6.4.
The variety $\widetilde{\mathcal{M}}_{H}$ is a point iff $\mu\in W\lambda$.
Proof.
Assume $\mu\in W\lambda$. For $\theta>0$, $\widetilde{\mathcal{M}}_{H}$ corresponds to the $\mu$ weight space of the irreducible representation with highest weight $\lambda$. Hence it is nonempty. Since $\dim\widetilde{\mathcal{M}}_{H}=(\lambda,\lambda)-(\mu,\mu)=0$, $\widetilde{\mathcal{M}}_{H}$ is a point. It is known, see for example Section 2.1 of [3], that the affinization of $\widetilde{\mathcal{M}}_{H}$ is independent of generic $\theta$. Hence $\widetilde{\mathcal{M}}_{H}$ is a point for any $\theta$.
Now assume that $\widetilde{\mathcal{M}}_{H}$ is a point. Then $(\lambda,\lambda)=(\mu,\mu)$. Furthermore, $\mu$ must be a weight of the irreducible representation $V(\lambda)$
It follows that $\mu^{+}\leqslant\lambda$, where $\mu^{+}$ is the dominant representative in $W\mu$. Our goal is to check that $\mu^{+}=\lambda$. Indeed note that
$0=(\lambda,\lambda)-(\mu^{+},\mu^{+})=(\lambda,\lambda-\mu^{+})+(\mu^{+},\lambda-\mu^{+})=(\lambda-\mu^{+},\lambda-\mu^{+})+2(\mu^{+},\lambda-\mu^{+})$
and $(\lambda-\mu^{+},\lambda-\mu^{+})\geqslant 0$, $(\mu^{+},\lambda-\mu^{+})\geqslant 0$. So the only possibility for this sum to be zero is if $(\lambda-\mu^{+},\lambda-\mu^{+})=0$, hence $\lambda=\mu^{+}$. ∎
Remark 6.5.
One can show that for an arbitrary quiver $Q$ without loops the corresponding quiver variety $\widetilde{\mathcal{M}}_{H}$ is a point iff $\mu\in W\lambda$. The argument is standard. Let us provide a sketch. The representation $V(\lambda)$ is integrable. Now, to every pair $\eta$, $\alpha_{i}$ of a weight $\eta$ of $V(\lambda)$ and a simple root $\alpha_{i}$ we consider the set of weights $S_{\eta,\alpha_{i}}$ of weights of $V(\lambda)$ of the form $\eta+k\alpha_{i}$ for $k\in\mathbb{Z}$. It is a standard fact (see [21, Proposition 3.6(b)]) that $\exists p,q,\in\mathbb{Z}_{\geq 0}$ such that $S_{\eta,\alpha_{i}}=\{\eta+k\alpha_{i}\,\mid\,-p\leq k\leq q\}$ and $s_{\alpha_{i}}(\eta-p\alpha_{i})=\eta+q\alpha_{i}$. Note now that $(\eta+k\alpha_{i},\eta+k\alpha_{i})=2k^{2}+2k(\eta,\alpha_{i})+(\eta,\eta)$ is a strictly convex function of $k$. So it must attain its maxima exactly at p-p and at $q$. Using this observation, one can prove that if $\mu$ is a weight of $V(\lambda)$, then $(\mu,\mu)\leqslant(\lambda,\lambda)$ and the equality holds iff $\mu\in W\lambda$ (argue by the induction on $\langle\rho,\mu\rangle$, where $\rho$ is the sum of fundamental weights for $\mathfrak{g}_{Q}$).
Proposition $6.$
Conjectures 5.2 and 5.3 hold for ADE quivers.
Proof.
Conjecture 5.2 follows from Lemmas 6.3, and 6.4. To prove Conjecture 5.3, recall that (18) identifies the partial resolution $\widetilde{\mathcal{M}}_{C,\nu}$ with the convolution diagram $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$. Now, Conjecture 5.3 follows from [25, Proposition 5.7] which states that the multiplication morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is a $\mathsf{A}_{Q}$-equivariant open embedding. ∎
Let us finally deal with Conjecture 1.5.
Proposition 6.7.
Conjecture 1.5 holds for ADE quivers.
Proof.
The Coulomb branches we must consider are $\mathcal{M}_{C}=\overline{\mathcal{W}}^{\lambda}_{\mu}$ for $\mu\in W\lambda$. Recall the point $z^{\mu}$ lies in $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$. Consider the repellent $R^{\lambda}_{\mu}$ (see [19, Definition 1.8.3]) with respect to the cocharacter $2\rho^{\vee}\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}_{Q}$, given by the sum of the positive coroots, inside the smooth locus $\mathcal{W}^{\lambda}_{\mu}$. At the level of $\mathbb{C}$-points we have:
$R^{\lambda}_{\mu}=\{x\in\mathcal{W}^{\lambda}_{\mu}\,|\,\underset{t\rightarrow 0}{\operatorname{lim}}\,2\rho^{\vee}(t)\cdot x=z^{\mu}\}.$
Note that $R^{\lambda}_{\mu}$ is nonsingular (see [19, Proposition 1.7.6], or, alternatively, use Luna’s slice theorem) and $T_{z^{\mu}}R^{\lambda}_{\mu}=(T_{z^{\mu}}\overline{\mathcal{W}}^{\lambda}_{\mu})^{-}$, where - corresponds to taking the direct sum in $T_{z^{\mu}}\mathcal{W}^{\lambda}_{\mu}$ of negative $2\rho^{\vee}(\mathbb{C}^{\times})$-weight subspaces. Using the symplectic form on $\mathcal{W}^{\lambda}_{\mu}$ we obtain the isomorphism of representations of ${\mathsf{T}}_{Q}$:
$T_{z^{\mu}}\overline{\mathcal{W}}^{\lambda}_{\mu}\simeq(T_{z^{\mu}}R^{\lambda}_{\mu})\oplus\hbar(T_{z^{\mu}}R^{\lambda}_{\mu})^{\vee}$ (20)
It remains to describe the ${\mathsf{T}}_{Q}$-character of $T_{z^{\mu}}R^{\lambda}_{\mu}$. It follows from [24, Theorem 3.1(1)] that $R^{\lambda}_{\mu}$ is isomorphic to the repellent to $z^{\mu}\in\operatorname{Gr}^{\lambda}$.
So, $T_{z^{\mu}}R^{\lambda}_{\mu}$ is isomorphic to $(T_{z^{\mu}}\operatorname{Gr}^{\lambda})^{-}$. We have
$T_{z^{\mu}}\operatorname{Gr}^{\lambda}=T_{z^{\mu}}(G_{\mathcal{O}}\cdot z^{\mu})=\\ =T_{1}(G_{\mathcal{O}}/(G_{\mathcal{O}}\cap z^{\mu}G_{\mathcal{O}}z^{-\mu}))\simeq\bigoplus_{\alpha\in\Phi,\,k=0,1,\ldots,\langle\alpha,\mu\rangle-1}z^{k}\mathfrak{g}_{\alpha}.$ (21)
The ${\mathsf{A}}_{Q}$-weight of $z^{k}\mathfrak{g}_{\alpha}$ is $\alpha$ and the $\mathbb{C}^{\times}$-weight of $z^{k}\mathfrak{g}_{\alpha}$ is k-k. Note now that passing from $T_{z^{\mu}}\operatorname{Gr}^{\lambda}$ to $(T_{z^{\mu}}\operatorname{Gr}^{\lambda})^{-}$ corresponds to restricting to $\alpha\in\Phi^{-}_{\mu}$ in the sum (21). We conclude that
$T_{z^{\mu}}R^{\lambda}_{\mu}=(T_{z^{\mu}}\operatorname{Gr}^{\lambda})^{-}=\sum_{\alpha\in\Phi^{-}_{\mu}}\sum_{i=1}^{\langle\alpha,\mu\rangle}\hbar^{1-i}e^{\alpha}$
and the claim follows from (20).
Remark 6.8.
Note that for $\lambda$ being minuscule and $\mu\in W\lambda$ Proposition 6.7 already follows from [25, Proposition 4.19], so the main new point of Proposition 6.7 is that the same formula holds for arbitrary dominant $\lambda$ and $\mu\in W\lambda$. It is worth noting that in [25, Section 4] a much stronger statement is proven for $\lambda$ minuscule and $\mu\in W\lambda$. Namely, it is proven that $\overline{\mathcal{W}}^{\lambda}_{\mu}=\mathcal{W}^{\lambda}_{\mu}\simeq T^{*}\mathbb{A}^{n}$ as symplectic varieties and the action of ${\mathsf{T}}_{Q}$ is described explicitly on the right hand side of the isomorphism. The same argument works for arbitrary $\lambda$ and $\mu\in W\lambda$ being almost dominant (i.e., $\langle\alpha,\mu\rangle\geqslant-1$ for any positive root $\alpha$).
Combining Propositions 6.6 and 6.7 we get the formula for the tangent space to an arbitrary nonsingular $\mathsf{T}_{Q}$-fixed point $z^{\underline{\mu}}$ of a partial resolution $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ (recall that these points are in bijection with all possible decompositions $\mu=\mu_{1}+\ldots+\mu_{N}$ such that $\mu_{i}\in W\lambda_{i}$). This is a generalization of [25, Equation (5.8)]:
$T_{z^{\underline{\mu}}}\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}=\sum_{k=1}^{N}\sum_{\alpha\in\Phi^{-}_{\mu_{k}}}\sum_{i=1}^{\langle\alpha,\mu_{k}\rangle}\left(\hbar^{1-i-\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{\alpha}+\hbar^{i+\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{-\alpha}\right).$
6.3 About the restriction to real roots
Let us finish this section with a few remarks. When $Q$ is of type ADE, it is known that for $\mu$ being dominant $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is indeed a slice to the $G_{\mathcal{O}}$-orbit $\operatorname{Gr}^{\mu}$ of $z^{\mu}$ inside the closure of the $G_{\mathcal{O}}$-orbit of $z^{\lambda}$. This is not the case for arbitrary $\mu$ (that is the reason why $\overline{\mathcal{W}}^{\lambda}_{\mu}$ are called generalized slices), although by [24, Theorem 3.1(1)] it is always true that the repellent to $z^{\mu}\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ under the $\mathbb{C}^{\times}$-action coming from $2\rho^{\vee}$ coincides with the repellent to $z^{\mu}\in\overline{\operatorname{Gr}}^{\lambda}$.
Now assume that we are in the setting of Conjecture 1.5 (i.e., variety $\widetilde{\mathcal{M}}_{H}$ is a point), and we now allow $Q$ to be an arbitrary quiver without loops. This is equivalent to $\mu\in W\lambda$ (see Remark 6.5 above).
Assuming that the $2\rho^{\vee}$-fixed point $p\in\mathcal{M}_{C}$ is nonsingular and denoting by $R^{\lambda}_{\mu}\subset\mathcal{M}_{C}$ the $2\rho^{\vee}$-repellent to $p$ we see that Conjecture 1.5 is equivalent to the isomorphism of ${\mathsf{T}}_{Q}$-representations:
$T_{p}R^{\lambda}_{\mu}=\bigoplus_{\alpha\in\Phi^{-,\mathrm{re}}}\mathfrak{g}_{\alpha}[[z]]/\Big(\bigoplus_{\alpha\in\Phi^{-,\mathrm{re}}}\mathfrak{g}_{\alpha}[[z]]\cap z^{\mu}\bigoplus_{\alpha\in\Phi^{-,\mathrm{re}}}\mathfrak{g}_{\alpha}[[z]]z^{-\mu}\Big),$ (22)
where $\Phi^{-,\text{re}}$ are all real negative roots for $\mathfrak{g}=\mathfrak{g}_{Q}$. We do not know if the right hand side of (22) has a “geometric” meaning similar to the one in finite dimensional situation (when it is the tangent space to the repellent to $z^{\mu}\in\operatorname{Gr}_{G}^{\lambda}$). It is also not clear to us why only real roots appear in (22).
The following lemma sheds some light on the importance of real roots, confirming that the dimensions of left and right hand sides of (22) are indeed equal. Recall that $\rho$ is the sum of fundamental weights for $\mathfrak{g}=\mathfrak{g}_{Q}$.
Lemma 6.9.
Let $\lambda$ be a dominant coweight of $\mathfrak{g}$ and let $\mu\in W\lambda$. Then
$\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\langle\alpha,\mu\rangle=\langle\lambda-\mu,\rho\rangle.$
Proof.
Recall that $\mu\in W\lambda$ and pick $w\in W$ of minimal length such that $w(\lambda)=\mu$. It is a standard fact that $w$
has the following property: if $\beta$ is any positive root such that $\langle\beta,\lambda\rangle=0$, then $w\beta$ is also positive.
Now, we can write:
$\langle\rho,\lambda-\mu\rangle=\langle\rho,w^{-1}\mu-\mu\rangle=\langle w(\rho),\mu\rangle-\langle\rho,\mu\rangle=\langle w(\rho)-\rho,\mu\rangle.$
Note that $w(\rho)-\rho$ is equal to the sum of all real negative roots $\alpha$ such that $w^{-1}(\alpha)$ is positive. Also note that $\langle\alpha,\mu\rangle=\langle w^{-1}\alpha,\lambda\rangle$. So, it remains to check that for a negative real root $\alpha$, $w^{-1}(\alpha)$ is positive iff $\langle w^{-1}\alpha,\lambda\rangle>0$.
Assume first that $\langle w^{-1}\alpha,\lambda\rangle>0$. The coweight $\lambda$ is dominant, so $w^{-1}(\alpha)$ must be positive.
Now assume $\beta:=w^{-1}(\alpha)$ is positive. Since $\lambda$ is dominant, $\langle\beta,\lambda\rangle\geqslant 0$. If $\langle\beta,\lambda\rangle=0$, then $\beta$ is a positive root such that $w(\beta)=\alpha$ is negative. This contradicts the property of $w$. ∎
Remark 6.10.
From the proof of Lemma 6.9, we see that the key spot that accounts for the restriction to the real roots is the equality:
$w(\rho)-\rho=\sum_{\begin{subarray}{c}\alpha\in\Phi^{-,\mathrm{re}}\\ w^{-1}(\alpha)\in\Phi^{+}\end{subarray}}\alpha.$
7 Slant sums of Coulomb branches
In this section, we study the slant sum construction from the Coulomb branch perspective.
7.1 Integrable systems for Coulomb branches
Coulomb branches come equipped with a certain additional structure called an integrable system. Recall that
$\mathcal{M}_{C}=\operatorname{Spec}H_{*}^{G_{\mathcal{O}}}(\mathcal{R}),~\mathcal{A}_{\hbar}=H_{*}^{G_{\mathcal{O}}\rtimes\mathbb{C}^{\times}}(\mathcal{R}).$
We have a natural (left) action of the algebra $H^{*}_{G_{\mathcal{O}}}(\operatorname{pt})=H^{*}_{G}(\operatorname{pt})$ on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ as well as the action of $H^{*}_{G_{\mathcal{O}}\rtimes\mathbb{C}^{\times}}(\operatorname{pt})=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})$ on $\mathcal{A}$. Acting on identity, we obtain homomorphisms:
$H^{*}_{G}(\operatorname{pt})\rightarrow\mathbb{C}[\mathcal{M}_{C}],~H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})\rightarrow\mathcal{A}_{\hbar}$
defining integrable systems on $\mathcal{M}_{C}$ and $\mathcal{A}_{\hbar}$. In this way, we obtain (Poisson) commutative subalgebras of $\mathbb{C}[\mathcal{M}_{C}]$ and $\mathcal{A}$. The subalgebra $H:=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})|_{\hbar=1}\subset\mathcal{A}$ is called the Cartan subalgebra in [8]. It is also sometimes called Gelfand–Tsetlin subalgebra (see, for example, [42]). We can naturally identify it with the polynomial ring $\mathbb{C}[c_{i,k}\,|\,i\in Q_{0},\,k=1,\ldots,v_{i}]$, where $c_{i,k}=c_{i}^{G\times\mathbb{C}^{\times}}(V_{k})$ ($V_{k}$ is considered as a $G\times\mathbb{C}^{\times}$-equivariant bundle on $\operatorname{pt}$).
One reason why the algebra $H$ is important is because it allows one to define the notion of a Gelfand–Tsetlin character of a module over $\mathcal{A}$. Namely, whenever $M$ is a finitely generated $\mathcal{A}$-module such that the action of $H$ on it decomposes $M$ into the direct sum of finite dimensional generalized eigenspaces (such modules are called Gelfand–Tsetlin-modules in [42]), we can define its Gelfand–Tsetlin character as follows.
For any collection ${\bf{S}}=(S_{i})_{i\in Q_{0}}$ such that $S_{i}$ is a set of $\mathsf{v}_{i}$ unordered complex numbers, we can define
$M_{\bf{S}}=\{x\in M\,|\,\forall i\in Q_{0},\,k=1,\ldots,v_{i}~\exists N>0~\text{s.t}~(c_{i,k}-e_{k}(S_{i}))^{N}=0\}.$
Then:
$\chi_{GT}(M):=\sum_{{\bf{S}}}z^{\bf{S}}\operatorname{dim}M_{\bf{S}},$
where we treat $z^{\bf{S}}$ as formal parameters. For an ADE quiver $Q$, the Gelfand-Tsetlin character is one way to package the Frenkel-Reshetikhin $q$-character of a module. For short, we will write GT-character whenever we refer to the Gelfand-Tsetlin character.
7.2 Slant sum of Coulomb branches with one dimensional framing
Let $Q^{(1)}$ and $Q^{(2)}$ be quivers and fix dimension and framing vectors $\mathsf{v}^{(r)},\mathsf{w}^{(r)}$ for $r\in\{1,2\}$. We now make the following assumption.
Assumption 7.1.
We have $\mathsf{w}^{(2)}_{\star_{2}}=1$ for some $\star_{2}\in Q^{(2)}_{0}$ and $\mathsf{w}^{(2)}_{i}=0$ otherwise.
Choose a vertex $\star_{1}\in Q_{0}^{(1)}$ such that $\mathsf{v}^{(1)}_{\star_{1}}=1$. As in Section 2.2, we can form the slant sum, which is the quiver gauge theory for the quiver $Q:=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$ with $\mathsf{v}=\mathsf{v}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathsf{v}^{(2)}$ and $\mathsf{w}=\mathsf{w}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathsf{w}^{(2)}$. Recall that $\mathsf{v}=\mathsf{v}^{(1)}\sqcup\mathsf{v}^{(2)}$ and $\mathsf{w}_{i}=\mathsf{w}_{i}^{(1)}$ for $i\in Q_{0}^{(1)}$, $\mathsf{w}_{i}=\mathsf{w}^{(2)}_{i}$ for $i\in Q_{0}^{(2)}\setminus\{\star_{2}\}$, and $\mathsf{w}_{\star_{2}}=0$.
As in (6), we have ${\bf{N}}^{(1)}$, ${\bf{N}}^{(2)}$, and $\bf{N}$} which are acted on by the gauge groups $G_{\mathsf{v}^{(1)}}$, $G_{\mathsf{v}^{(2)}}$, and $G_{\mathsf{v}}$ respectively. These give rise to the corresponding Coulomb branches $\mathcal{M}_{C}^{(1)}$, $\mathcal{M}_{C}^{(2)}$, and $\mathcal{M}_{C}$ as well as their quantizations $\mathcal{A}^{(1)}$, $\mathcal{A}^{(2)}$, and $\mathcal{A}$.
Let
$\mathsf{A}_{Q^{(1)}}=\pi_{1}(G_{\mathsf{v}^{(1)}})^{\wedge},~\mathsf{A}_{Q^{(2)}}=\pi_{1}(G_{\mathsf{v}^{(2)}})^{\wedge},~\mathsf{A}_{Q}=\pi_{1}(G_{\mathsf{v}})^{\wedge}$
be the tori acting naturally on $\mathcal{M}_{C}^{(1)}$, $\mathcal{M}_{C}^{(2)}$, and $\mathcal{M}_{C}$ respectively (see Section 5.3 above). We will denote by $\mathsf{T}_{Q^{(1)}}$, $\mathsf{T}_{Q^{(2)}}$, and $\mathsf{T}_{Q}$ their products with $\mathbb{C}^{\times}_{\hbar}$.
Consider the isomorphism
[…]
defined as follows. Recall that $\operatorname{GL}(V_{\star_{1}})=\mathbb{C}^{\times}$ and for $k\in Q^{(2)}_{0}$ let $\iota_{k}\colon\operatorname{GL}(V_{\star_{1}})\hookrightarrow\operatorname{GL}(V^{(2)}_{k})$ be the embedding given by $t\mapsto\operatorname{diag}(t,t,\ldots,t)$. Then $\varphi(g^{(1)},g^{(2)})_{k}$ is equal to $g_{k}$ for $k\in Q_{0}^{(1)}$, and is equal to $\iota_{k}(g_{\star_{1}})g^{(2)}_{k}$ for $k\in Q_{0}^{(2)}$.
Applying $\pi_{1}(\bullet)^{\wedge}$ to both sides of the isomorphism $\varphi$, we obtain the identification:
[…]
The isomorphism $\varphi$ also induces identifications:
$H^{*}_{G_{\mathsf{v}^{(1)}}}(\operatorname{pt})\otimes H^{*}_{G_{\mathsf{v}^{(2)}}}(\operatorname{pt})\simeq H^{*}_{G_{\mathsf{v}}}(\operatorname{pt})$
[…]
Proposition 7.2.
There exist $\mathsf{T}_{Q}$-equivariant isomorphisms of Poisson varieties and algebras:
$\mathcal{M}_{C}\simeq\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)},~\mathcal{A}\simeq\mathcal{A}^{(1)}\otimes\mathcal{A}^{(2)}.$
This isomorphism is compatible with the identification (23) of integrable systems.
Proof.
It follows from the definitions that $\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)}$ is the Coulomb branch for the pair $(G_{\mathsf{v}^{(1)}}\times G_{\mathsf{v}^{(2)}},{\bf{N}}^{(1)}\oplus{\bf{N}}^{(2)})$. So, to construct the desired isomorphism it is enough to construct isomorphisms
[…]
such that
$\psi((g^{(1)},g^{(2)})\cdot(n^{(1)},n^{(2)}))=\varphi(g^{(1)},g^{(2)})\cdot\psi(n^{(1)},n^{(2)})$~
for all $g^{(r)}\in G_{\mathsf{v}^{(r)}}$ and $n^{(r)}\in{\bf{N}}^{(r)}$, $r\in\{1,2\}$. We have already constructed the isomorphism $\varphi$.
[…]
It follows from the definitions that the maps $\varphi$, $\psi$ satisfy the desired properties. Compatibility with integrable systems and ${\mathsf{T}}_{Q}$-equivariance follows from construction. ∎
Remark 7.3.
Fixing arbitrary pair of characters $\theta^{(1)}\colon G_{\mathsf{v}^{(1)}}\rightarrow\mathbb{C}^{\times}$, $\theta^{(2)}\colon G_{\mathsf{v}^{(2)}}\rightarrow\mathbb{C}^{\times}$ and using the identification $\varphi$ we obtain the character $\theta\colon G_{\mathsf{v}}\rightarrow\mathbb{C}^{\times}$. The same argument as in the proof of Proposition 7.2 above shows that the identification $\mathcal{M}_{C}\simeq\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)}$ lifts to the $\mathsf{T}_{Q}$-equivariant identification of partially resolved Coulomb branches $\widetilde{\mathcal{M}}_{C,\theta}\simeq\widetilde{\mathcal{M}}_{C,\theta^{(1)}}^{(1)}\times\widetilde{\mathcal{M}}_{C,\theta^{(2)}}^{(2)}$. Similarly, if we have any pair of tori ${\mathsf{A}}^{(1)},\mathsf{A}^{(2)}$ acting on ${\bf{N}}^{(1)}$, ${\bf{N}}^{(2)}$, then, using the identification $\psi$, we obtain the action of ${\mathsf{A}}^{(1)}\times\mathsf{A}^{(2)}$ on $\bf{N}$} and then the isomorphism of Proposition 7.2 can be upgraded to deformations.
One immediate corollary of Proposition 7.2 is
Corollary 7.4.
If Conjecture 5.2 holds for $\mathcal{M}_{C}^{(1)}$, $\mathcal{M}_{C}^{(2)}$, then it also holds for $\mathcal{M}_{C}$. If $p^{(i)}$, $i=1,2$ are fixed points of $\mathcal{M}_{C}^{(i)}$, we denote by $p^{(1)}\#p^{(2)}$ the corresponding (unique) fixed point of $\mathcal{M}_{C}$.
Remark 7.5.
Proposition 7.2 allows one to extract the character of $T_{p^{(1)}\#p^{(2)}}\mathcal{M}_{C}$ from the characters of $T_{p^{(i)}}\mathcal{M}_{C}^{(i)}$, $i=1,2$. Proposition 7.2 also allows us to compute the GT-character for an irreducible module $L(p^{(1)}\#p^{(2)})$ over $\mathcal{A}$ corresponding to $p^{(1)}\#p^{(2)}$ (see Section 5.4 above) from the GT-characters of irreducible modules for $\mathcal{A}^{(i)}$ corresponding to $p^{(i)}$.
Let us now relax Assumption 7.1.
Assumption 7.6.
We have $\mathsf{w}^{(2)}_{\star_{2}}=1$ for some $\star_{2}\in Q^{(2)}_{0}$.
As above, we fix the identification $W^{(2)}_{\star_{2}}\simeq V^{(1)}_{\star_{1}}\simeq\mathbb{C}$. Identifying $\mathbb{C}^{\times}\simeq GL(W^{(2)}_{\star_{2}})$ we obtain the action $\mathbb{C}^{\times}\curvearrowright{\bf{N}}^{(2)}$. Consider the following embedding:
$G_{\mathsf{v}^{(1)}}\times G_{\mathsf{v}^{(2)}}\hookrightarrow G_{\mathsf{v}}\times\mathbb{C}^{\times},~(g^{(1)},g^{(2)})\mapsto(g^{(1)},g^{(2)},g^{(1)}_{\star_{1}})$. (24)
It follows from the definitions that this embedding intertwines the actions on ${\bf{N}}={\bf{N}}^{(1)}\oplus{\bf{N}}^{(2)}$. The cokernel of (24) identifies with $\mathbb{C}^{\times}$.
Applying [8, Proposition 3.18] we obtain the following proposition describing the Coulomb branch $\mathcal{M}_{C}$ as a Hamiltonian reduction of the product of Coulomb branches.
Proposition 7.7.
We have isomorphisms:
$\mathcal{M}_{C}\simeq(\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}(G_{\mathsf{v}^{(2)}}\times\mathbb{C}^{\times},{\bf{N}}^{(2)}))/\!\!/\mathbb{C}^{\times},~\mathcal{A}_{C}\simeq(\mathcal{A}_{C}^{(1)}\otimes\mathcal{A}_{C}(G_{\mathsf{v}^{(2)}}\times\mathbb{C}^{\times},{\bf{N}}^{(2)}))/\!\!/\mathbb{C}^{\times}.$